### v1

# Optimal Stopping with a Predicted Prior

Optimal Stopping with a Predicted Prior
Tian Bai [^1] University of Bergen. Email: tian.bai@uib.no. This work was done when the author was at the University of Hong Kong.    Zhiyi Huang [^2] The University of Hong Kong. Email: zhiyi@cs.hku.hk, dongchen.li@connect.hku.hk    Chui Shan Lee [^3] Hong Kong University of Science and Technology. Email: tracylcs@ust.hk. This work was done when the author was at the University of Hong Kong.    Dongchen Li22footnotemark: 2
(November 2025)
There are two major models of value uncertainty in the optimal stopping literature: the secretary model, which assumes no prior knowledge, and the prophet inequality model, which assumes full information about value distributions. In practice, decision makers often rely on machine-learned priors that may be erroneous. Motivated by this gap, we formulate the model of optimal stopping with a predicted prior to design algorithms that are both consistent, exploiting the prediction when accurate, and robust, retaining worst-case guarantees when it is not.
Existing secretary and prophet inequality algorithms are either pessimistic in consistency or not robust to misprediction. A randomized combination only interpolates their guarantees linearly. We show that a family of bi-criteria algorithms achieves improved consistency-robustness trade-offs, both for maximizing the expected accepted value and for maximizing the probability of accepting the maximum value. We further prove that for the latter objective, no algorithm can simultaneously match the best prophet inequality algorithm in consistency, and the best secretary algorithm in robustness.
1 Introduction
Optimal stopping problems study a fundamental dilemma in decision‑making: to seize an expiring opportunity now, or to wait for a better chance that may never come. A classic example is the secretary problem. In this problem, $n$ values are revealed to a decision maker in a random order. After observing each value, the decision maker immediately decides whether to accept the value. The decision maker can accept at most one value and aims to maximize its expectation. An optimal solution is Dynkin’s algorithm [dynkinOptimumChoiceInstant1963, fergusonWhoSolvedSecretary1989], which rejects the first $\lfloor\frac{n}{\mathrm{e}}\rfloor$ values, and then accepts the first value that exceeds all previous ones. It guarantees that the expected accepted value is at least $\frac{1}{\mathrm{e}}$ times the maximum. This ratio of the algorithm’s expected accepted value to the maximum value in hindsight is called the competitive ratio.
In sharp contrast to the minimal prior knowledge assumed in the secretary problem, prophet inequality considers values drawn independently from known prior distributions. In the i.i.d. case where values are drawn from a common distribution, the optimal algorithm guarantees that the expected accepted value is at least approximately $0.745$ times the expected maximum value [hillComparisonsStopRule1982, kertzStopRuleSupremum1986, correaPostedPriceMechanisms2017, liuVariableDecompositionProphet2021], substantially better than prior-free secretary guarantee of $\frac{1}{\mathrm{e}}\approx 0.368$.
On the one hand, the secretary problem does not capture the abundant data available to decision makers in this era. On the other hand, the assumption of complete knowledge of the prior in prophet inequality is idealized. What real-world decision makers often have are machine-learned and potentially erroneous predicted priors. duttingPostedPricingProphet2019 explored this direction for the general prophet inequality problem. They examined how mispredicted priors affect existing methods and showed a degradation in performance proportional to the prediction error. This can be viewed as an analysis of the smoothness of algorithms under the framework of algorithms with predictions [mitzenmacher2022algorithms].
1.1 Optimal Stopping with a Predicted Prior
We formulate the problem of optimal stopping with a predicted prior and focus on the trade-off between consistency and robustness, i.e., the performance when the predicted prior is correct and when it may be arbitrarily wrong. Consider the optimal stopping problem with $n$ values drawn i.i.d. from an unknown prior distribution $F$. The algorithm is given the number of values $n$ and a predicted prior $\tilde{F}$ in the beginning. Then, the algorithm decides whether to accept each value upon its arrival, based on the history of observed values and the predicted prior $\tilde{F}$. We consider two standard objectives: 1) maximizing the expected accepted value (MaxExp), which we have already discussed, and 2) maximizing the probability of accepting the maximum value (MaxProb).
An algorithm is $\alpha$-consistent if it is $\alpha$-competitive when the predicted prior is correct. It is $\beta$-robust if it remains $\beta$-competitive even when the (mis)predicted prior is arbitrary. Ideally, we want an algorithm to be both $\alpha$-consistent and $\beta$-robust with $\alpha$ and $\beta$ being as large as possible. Such an algorithm leverages prior knowledge when the predicted prior is correct, while retaining meaningful worst-case guarantees.
The $\frac{1}{\mathrm{e}}$-competitive Dynkin’s algorithm is $\frac{1}{\mathrm{e}}$-consistent and $\frac{1}{\mathrm{e}}$-robust for both objectives. The optimal $\alpha^{*}_{\textrm{MaxExp}}\approx 0.745$-competitive MaxExp algorithm by Hill and Kertz [hillComparisonsStopRule1982, kertzStopRuleSupremum1986] and $\alpha^{*}_{\textrm{MaxProb}}\approx 0.580$-competitive MaxProb algorithm by gilbertRecognizingMaximumSequence are $\alpha^{*}_{\textrm{MaxExp}}$ and $\alpha^{*}_{\textrm{MaxProb}}$-consistent respectively, but only 0-robust (see Section 2). Further, randomly running either Dynkin’s algorithm or an i.i.d. prophet inequality algorithm linearly interpolates between their consistency and robustness. Finally, the optimality of these competitive ratios implies that no algorithm is better than $\frac{1}{\mathrm{e}}$-robust, or better than $\alpha^{*}_{\textrm{MaxExp}}$-consistent for MaxExp, or better than $\alpha^{*}_{\textrm{MaxProb}}$-consistent for MaxProb. These baseline algorithms and impossibilities leave a significant gap between them.
1.2 Overview of Results
This paper introduces two new algorithms for the MaxExp and MaxProb objectives respectively, under a unified framework, and a hardness result for MaxProb. See Figure 1 for an illustration.
(a) MaxExp
(b) MaxProb
Figure 1: Illustration of our results and a comparison with the baselines
Bi-criteria Algorithms.
We define a family of bi-criteria algorithms parameterized by the predicted prior $\tilde{F}$ and a threshold function $\theta$. Upon observing each value and when no value has been accepted, a bi-criteria algorithm accepts the current value if:
  1. 1.
    It is best-so-far, i.e., greater than or equal to all previous values; (prior-free criterion)
  2. 2.
    Its cdf value w.r.t. $\tilde{F}$ exceeds a time-dependent threshold $\theta(t)$. (prior-dependent criterion)
For now, readers may interpret $t$ as the time step by which the value is observed. Later, we will also consider a continuous-time model to simplify the analysis (see Section 2 for a discussion).
We set the thresholds as follows. Before some time $\lambda_{1}$, the threshold is $1$, and the algorithm rejects all values. After some time $\lambda_{2}$, the threshold is 0, and the algorithm’s decisions depend solely on the prior-free best-so-far criterion. These two prior-free phases yield robustness. In fact, we derive this design from first principles, as we will explain shortly in Section 1.3.
We obtain new algorithmic results for both MaxExp and MaxProb under this framework.
Theorem 1 (MaxExp algorithm).
Suppose $0\leq\beta\leq\frac{1}{e}$ and $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$. Further, suppose there is a left-continuous non-increasing function $\theta:[\lambda_{1},\lambda_{2}]\to[0,1]$ satisfying:
$\theta(\lambda_{1})$ $~=~1~,$
$\int_{z}^{1}\int_{0}^{t}\,\frac{1}{t}\theta(\max\{s,z\})^{t}\,\mathrm{d}s\,\mathrm{d}t$ $~\geq~\alpha\cdot\theta(z)~,\qquad\forall\lambda_{1}\leq z\leq\lambda_{2}~.$
Then, there is an $\alpha$-consistent $\beta$-robust algorithm for the MaxExp objective.
To plot the curve in Figure 1(a), we rely on numerical methods to solve the differential equation in Theorem 1, which does not appear to admit a closed-form solution. See Appendix B.
Theorem 2 (MaxProb algorithm).
Suppose $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$. There is an $\alpha$-consistent $\beta$-robust algorithm for the MaxProb objective for any $\beta\in[0,\frac{1}{e}]$ and
$\alpha=\beta+\int_{\lambda_{1}}^{\lambda_{2}}\int_{s}^{1}\,\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~,$
where $\gamma\approx 0.804$ is the solution to $\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!k}=1$.
We further prove that certain consistency-robustness trade-offs cannot be achieved for the MaxProb objective. In particular, there is no “best-of-both-world” algorithm.
Theorem 3 (MaxProb hardness, informal version of Theorem 15).
There is an $\alpha$-consistent $\beta$-robust algorithm for the MaxProb objective only if a polytope $\mathcal{P}(\alpha,\beta)$ is feasible. In particular, no algorithm is both $\alpha^{*}_{\textrm{MaxProb}}$-consistent and $\frac{1}{\mathrm{e}}$-robust.
1.3 Overview of Techniques
From Mispredicted Prior to Misspecified Thresholds.
A key property of bi-criteria algorithms is that they provide two equivalent viewpoints for robustness. By definition, we may consider a real prior $F$ and a (mis)predicted prior $\tilde{F}$ that misleads the algorithm. Equivalently, we may also consider an algorithm with the correct prior $F$ but a misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta$, since $\tilde{F}(x)\geq\theta(t)$ if and only if $F(x)\geq(F\circ\tilde{F}^{-1}\circ\theta)(t)$. The latter viewpoint is much easier to work with — we just need to restrict the misspecified threshold $F\circ\tilde{F}^{-1}\circ\theta(t)$ via our choice of threshold $\theta$. Further, the only restrictions are that whenever $\theta(t)$ is 0 or $1$, $F\circ\tilde{F}^{-1}\circ\theta(t)$ has the same value of 0 or $1$. This observation leads to our design of thresholds with three phases: 1) a prior-free phase before time $\lambda_{1}$ where $\theta(t)=1$ and the algorithm rejects all values; 2) a prior-dependent phase from time $\lambda_{1}$ to $\lambda_{2}$ where $\theta(t)$ is between 0 and $1$; and 3) a final prior-free phase after time $\lambda_{2}$ where $\theta(t)=0$ and the algorithm takes any value that is best-so-far. See Section 3 for a detailed discussion, and Section 5 for the robustness analysis.
Implicit Sharding.
We show that optimal stopping with a predicted prior is hardest when the number of values tends to infinity, using an implicit sharding technique that may be of independent interest. This is inspired by the sharding technique by harbNewProphetInequalities2025, which breaks each value into many i.i.d. shards based on the prior, such that the original value equals the maximum shard. However, the sharding technique requires full knowledge of the real prior, and thus, does not apply to optimal stopping with a predicted prior. In contrast, implicit sharding directly works with the original values, and only uses the shards implicitly in the analysis through a coupling argument. See Section 4 and specifically Theorem 7 for the formal argument.
Tractable Factor-revealing Linear Programs.
We prove the hardness result by constructing factor-revealing linear programs whose optimal values upper-bound the achievable consistency-robustness trade-offs. Exponential-size linear programs can easily capture all possible algorithms, which map possible sequences of observed values to binary acceptance decisions. The main challenge is to reduce the size without losing much characterization power.
We compress the algorithms’ state space by restricting the real prior to be a conditional distribution of the predicted prior, subject to an upper bound $k$ on the values. Intuitively, the only information an algorithm may conclude from the observed values $x_{1},\ldots,x_{t}$ is $k\geq\max_{1\leq i\leq t}x_{i}$ under this setup. We formalize this intuition, showing that we may, without loss of generality, consider minor-oblivious algorithms whose decisions depend solely on the maximum observed value, not on the detailed observed value sequence. This reduces the state space from an exponential number of possible value sequences to a polynomial number of possible maximum values. See Section 8 for the factor-revealing linear programs and the hardness.
1.4 Related Work
Prophet Inequalities with Full Prior Information.
hillComparisonsStopRule1982 proposed a $(1-\frac{1}{\mathrm{e}})$-competitive algorithm for i.i.d. prophet inequality and the MaxExp objective. This was improved to $0.738$ by abolhassaniBeating11Ordered2017, and finally to approximately $0.745$ by correaPostedPriceMechanisms2017. The latter ratio is tight, matching a hardness result by Hill and Kertz [hillComparisonsStopRule1982, kertzStopRuleSupremum1986]. For general prophet inequality with non-identical distributions, the optimal competitive ratio is $0.5$ [krengel1978semiamarts, samuel1984comparison]. The prophet secretary problem [esfandiariProphetSecretary2015] interpolates between general and i.i.d. prophet inequalities, with values drawn from non-identical distributions and arriving by a random order. The optimal competitive ratio for prophet secretary is open, with the best lower and upper bounds being $0.688$ [chenProphetSecretaryMatching2025] and $0.723$ [giambartolomeiProphetInequalitiesSeparating2023].
For the MaxProb objective, gilbertRecognizingMaximumSequence gave an optimal algorithm for the i.i.d. case, while samuels1982 derived an explicit expression of the optimal competitive ratio of approximately $0.580$. For non-identical distributions, esfandiariProphetsSecretariesMaximizing2020 obtained an optimal $\frac{1}{\mathrm{e}}$ competitive ratio in the prophet inequality model, and recovered the $0.580$ ratio in the prophet secretary model under a “no-superstar” assumption. nutiSecretaryProblemDistributions2022 later removed this assumption.
See correaRecentDevelopmentsProphet2019 for a survey. We stress that none of these prophet inequality algorithms provides non-trivial robustness.
Prophet Inequalities with Partial Prior Information.
There has been extensive research on the sample-based model, in which the algorithm has access to i.i.d. samples from the prior. One line of research focuses on getting constant-factor prophet inequalities using one (or fewer) sample from each distribution [azar2014prophet, rubinsteinOptimalSinglechoiceProphet2020, gravinOptimalProphetInequality2022, ezraProphetInequalitySamples2024]. Notably, the optimal $\frac{1}{2}$-competitive prophet inequality can be achieved with only one sample from each distribution [rubinsteinOptimalSinglechoiceProphet2020, gravinOptimalProphetInequality2022]. Another line of work studies the sample complexity of learning a nearly optimal algorithm [diakonikolasLearningOnlineAlgorithms2021, correaUnknownIIDProphets2021, correaTwosidedGameGoogol2020, rubinsteinOptimalSinglechoiceProphet2020, cristiProphetInequalitiesRequire2023, guo2021generalizing, correaSampledrivenOptimalStopping2021, jin2024sample], culminating in the state-of-the-art $\tilde{\mathcal{O}}(\frac{n}{\varepsilon})$ sample complexity for i.i.d. prophet inequality [correaSampledrivenOptimalStopping2021], and $\tilde{\mathcal{O}}(\frac{1}{\varepsilon^{2}})$ sample complexity per distribution for the general case [jin2024sample]. cristiProphetInequalitiesRequire2023 gave a unified proof that $\mathcal{O}(\mathrm{poly}(\frac{1}{\epsilon}))$ samples per distribution suffice for several major variants of prophet inequalities.
Other models of partial information include learning nearly optimal algorithms via quantile queries [liProphetInequalityIID2023], bandit feedback [gatmiry2024bandit], or simple posted-pricing policies [perezsalazarIIDProphetInequality2023]. These models and the sample-based model are incomparable to ours. They assume partial information from the correct prior, while the model of this paper considers a predicted prior that may be arbitrarily wrong.
Algorithms with Predictions.
Our work falls within the algorithms with predictions framework [mitzenmacher2022algorithms] for augmenting worst-case analysis of algorithms with machine-learned advice. Consistency, robustness, and smoothness are the primary performance metrics, capturing the algorithms’ performance when the prediction is correct, completely wrong, and partially wrong. The framework has been applied to many optimization problems, including scheduling [kumar2018improving] and caching [lykouris2021competitive]. In the context of optimal stopping, antoniadisSecretaryOnlineMatching2020 studied a prediction of the maximum value, fujiiSecretaryProblemPredictions2024 assumed having a predicted sequence of values, and braun2024secretary considered a prediction of the gap between the maximum value and the $k$-th largest. In contrast, this paper considers distributional predictions, which aligns with the literature of prophet inequality and the machine-learned priors in practice. As mentioned earlier, duttingPostedPricingProphet2019 also considered mispredicted prior with a focus on smoothness, while this paper examines the trade-off between consistency and robustness.
2 Preliminaries
Let $\mathbb{N}$ denote the set of positive integers. For any $n\in\mathbb{N}$, define $n]\coloneqq\{1,2,\ldots,n\}$. Let $\mathcal{D}$ be the set of all probability distributions supported on $0,\infty)$. We adopt the standard abbreviations cdf, pdf, and pmf for cumulative distribution function, probability density function, and probability mass function. We often use the same symbol for a distribution and its cdf (e.g., $F$), when clear from context. Unless stated otherwise, expectations and probabilities are taken over the randomness of the values and any internal randomness of the algorithm.
2.1 Optimal Stopping
Consider $n$ values $x_{1},x_{2},\ldots,x_{n}$ drawn i.i.d. from a prior distribution $F\in\mathcal{D}$. The values arrive one by one. Upon the arrival of each value $x_{i}$, the algorithm observes the realized value, and then must immediately decide whether to accept $x_{i}$ and stop, or to reject it and continue. At most one value may be accepted.
Fix any algorithm, let Alg be the index of the value it accepts, and thus, $x_{\textsc{Alg}}$ be the corresponding accepted value. We define $x_{\textsc{Alg}}=0$ if the algorithm rejects all values.
We consider the maximization problem, where larger values are preferable. The literature has studied two objectives. Let $x_{\textsc{Opt}}$ denote the maximum value among all $n$ values, i.e., $x_{\textsc{Opt}}\coloneqq\max_{i\in[n]}x_{i}$. The max expectation (MaxExp) objective considers the expected value of the algorithm’s accepted decision, $\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]$. The corresponding competitive ratio is
$\frac{\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]}{\operatorname{\mathbf{E}}[x_{\textsc{Opt}}]}~.$
The max probability (MaxProb) objective considers the probability of accepting the maximum:
$\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]~.$
This coincides with the competitive ratio since the offline optimal choice is always maximum. An algorithm is $\Gamma$-competitive if its competitive ratio is at least $\Gamma$.
2.2 Optimal Stopping with a Predicted Prior
We assume that the number of values $n$ is known. Depending on what the algorithm knows about the prior distribution $F$, we have different optimal stopping problems. If nothing is known, we have essentially the secretary problem [dynkinOptimumChoiceInstant1963, fergusonWhoSolvedSecretary1989]. [^4] The classical secretary problem considers random order without prior. Nonetheless, correaProphetInequalitiesIID2019 showed that the optimal competitive ratio is still $\frac{1}{\mathrm{e}}$ for an unknown prior. With complete information of $F$, we have the i.i.d. prophet inequality [correaUnknownIIDProphets2021].
This paper considers a new model between the two extremes, which we dub optimal stopping with a predicted prior. In this model, the algorithm is given a possibly erroneous predicted prior $\tilde{F}$. Hence, upon receiving each value $x_{i}$, the algorithm decides whether to accept based on the number of values $n$, the predicted prior $\tilde{F}$, and the realizations of the values observed thus far including $x_{i}$.
Following the literature of algorithms with predictions [mitzenmacher2022algorithms], we consider the consistency and robustness of stopping algorithms.
Definition 1 (Consistency).
An algorithm is $\alpha$-consistent if for any prior distribution $F\in\mathcal{D}$, it is $\alpha$-competitive when the predicted prior is accurate, i.e., when $\tilde{F}=F$.
Definition 2 (Robustness).
An algorithm is $\beta$-robust if for any prior distribution $F\in\mathcal{D}$, it is $\beta$-competitive regardless of which predicted prior $\tilde{F}$ is given.
2.3 Simplifying Assumptions for Algorithm Analyses
For our algorithmic results, we adopt two standard assumptions from the literature that simplify the analysis and are without loss of generality.
Continuous-time Random-order Model.
Besides the above classical model, which we will refer to as the discrete-time model, we may artificially introduce a random order of arrivals by letting each value $x_{i}$ with $i\in[n]$ arrive in a continuous time horizon $[0,1]$, drawing its arrival time $t_{i}$ independently and uniformly from $[0,1]$. Since the values are drawn i.i.d. from $F$, the problem remains the same.
An algorithm for the discrete-time model also works in the continuous-time model by ignoring the arrival time. The converse is also true. We can simulate the continuous-time model by first drawing $n$ arrival times independently and uniformly from $[0,1]$, and then mapping these arrival times in ascending order to the values in their arrival order.
Full-support Continuous Priors.
For the analysis, we further assume that both $F$ and $\tilde{F}$ are continuous with full support on $0,\infty)$. This assumption is without loss of generality, e.g., by independently perturbing each value with non-negative noise drawn from an exponential distribution with parameter $\varepsilon\to 0$; see correaProphetSecretaryBlind2021 for a similar treatment.
For any full-support continuous distribution $F\in\mathcal{D}$, define its right‑continuous inverse as:
$F^{-1}(q)\coloneqq\inf\big\{x\geq 0:F(x)\geq q\big\},\quad\forall q\in[0,1].$
Importantly, $F$ and $F^{-1}$ are strictly increasing on the interiors of their domains. Further, we have the boundary conditions $F^{-1}(0)=0$ and $F(0)=0$, and by convention $F^{-1}(1)=\infty$ and $F(\infty)=1$.
3 Bi-criteria Algorithms
We start with two definitions that will be repeatedly used throughout the paper.
Definition 3.
The prefix maximum for any time $t\in[0,1]$ is the maximum value arrived before time $t$, i.e., $\max_{i\,:\,t_{i}<t}x_{i}$, with the convention that it equals 0 if no value arrives before $t$.
Definition 4.
A value $x$ arriving at time $t$ is best-so-far if $x$ is greater than or equal to the prefix maximum for time $t$.
We consider the following family of algorithms parameterized by the predicted prior $\tilde{F}$ and a threshold function $\theta$. Throughout the paper, threshold functions are always non-increasing maps from $[0,1]$ to $[0,1]$.
Bi-criteria Algorithm $\mathcal{A}(\tilde{F},\theta)$ For each value $x_{i}$ arriving at time $t_{i}$ when no value has been accepted, accept $x_{i}$ if: 1. Value $x_{i}$ is best-so-far; and 2. The cdf value of $x_{i}$ w.r.t. $\tilde{F}$ is greater than the threshold at time $t_{i}$, i.e., if $\tilde{F}(x_{i})>\theta(t_{i})$.
The first prior-free criterion is without loss of optimality for MaxProb and is critical for the robustness guarantees for both MaxProb and MaxExp. The second prior-dependent criterion represents how the algorithm uses the predicted prior to guide its decisions.
Specifically, if $\theta(t)=1$, the algorithm accepts values with probability 0 at time $t$, regardless of how large the values are. If $\theta(t)=0$, the algorithm accepts any value that is best-so-far at time $t$, regardless of the value’s relative standing w.r.t. the predicted prior.
Algorithms under this framework satisfy many properties, including a simple necessary and sufficient condition for the algorithm to reject all values before time $t$, as follows. As discussed in Section 2.3, we assume the distribution is continuous and thus ties occur with probability zero.
Proposition 4.
The algorithm $\mathcal{A}$ rejects all values before time $t$ if and only if the prefix maximum $y$ for time $t$ and the arrival time $s$ of $y$ satisfy $\tilde{F}(y)\leq\theta(s)$.
Proof.
If $\tilde{F}(y)\leq\theta(s)$, then all values $x$ arriving before time $s$ are rejected because the cdf value is below the threshold, as $\tilde{F}(x)\leq\tilde{F}(y)\leq\theta(s)\leq\theta(t)$ for any $t\leq s$. Values arriving between $s$ and $t$ are also rejected because they do not exceed $y$ and therefore cannot be best-so-far.
If $\tilde{F}(y)>\theta(s)$, then $y$ is best-so-far and has cdf value above the threshold. Hence, the algorithm accepts either $y$ or a value before $y$. ∎
Corollary 5.
Suppose a value $x$ arrives at time $t$, and its prefix maximum $y$ arrives at time $s<t$. The algorithm accepts $x$ if and only if 1) $\tilde{F}(x)>\theta(t)$, 2) $x\geq y$, and 3) $\tilde{F}(y)\leq\theta(s)$.
3.1 Examples
This family encompasses many classical algorithms as special cases. We present three examples.
Example 1 (Dynkin’s algorithm [dynkinOptimumChoiceInstant1963], continuous-time variant).
Dynkin’s algorithm accepts the first best-so-far value after rejecting the first $\lfloor\frac{n}{\mathrm{e}}\rfloor$ values. The continuous-time variant (e.g., feldmanImprovedCompetitiveRatios2011) accepts the first best-so-far value after time $\frac{1}{\mathrm{e}}$, which corresponds to the threshold function $\theta(t)=1$ for $t\in[0,\frac{1}{\mathrm{e}}]$, and $\theta(t)=0$ for $t\in(\frac{1}{\mathrm{e}},1]$.
Dynkin’s algorithm is $\frac{1}{\mathrm{e}}$-competitive for both MaxProb and MaxExp, without any information about the prior [dynkinOptimumChoiceInstant1963, feldmanImprovedCompetitiveRatios2011]. Hence, it is $\frac{1}{\mathrm{e}}$-consistent and $\frac{1}{\mathrm{e}}$-robust.
Example 2 (Gilbert-Mosteller algorithm [gilbertRecognizingMaximumSequence], continuous-time variant).
Gilbert and Mosteller calculated the optimal thresholds for MaxProb through backward induction with a correct prior, i.e., $\tilde{F}=F$. Their algorithm achieves a competitive ratio of approximately $0.580$, with the following analytical form due to samuels1982: [^5] The competitive ratio presented in samuels1982 is $\mathrm{e}^{-\gamma}+(\mathrm{e}^{\gamma}-\gamma-1)\int_{1}^{\infty}\frac{\mathrm{e}^{-\gamma x}}{x}\,\mathrm{d}x$ (fixing a typographical negative sign before the second term), which equals our expression through transformations.
$\int_{0}^{1}\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~,$
where $\gamma\approx 0.804$ is the solution to $\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!k}=1$.
The continuous-time variant under our framework corresponds to the threshold function given by the following equation:
$\int_{t}^{1}\frac{(1-s+s\theta(t))^{n-1}-\theta(t)^{n-1}}{1-s}\,\mathrm{d}s=\theta(t)^{n-1}~.$
It achieves the same consistency $0.580$ for MaxProb as Gilbert-Mosteller algorithm (as a special case of Theorem 2). However, the algorithm is merely 0-robust. To illustrate, consider the case where the prior $F$ is uniform over $[0,1]$, whereas the predicted prior $\tilde{F}$ is uniform over $[2,3]$. In this case, the algorithm rejects all values, as they have cdf value 0 w.r.t. the predicted prior.
Example 3 (Single-threshold algorithm [ehsaniProphetSecretaryCombinatorial2018], i.i.d. case).
Let $\theta(t)=1-\tfrac{1}{n}$ for $t\in[0,1]$. [^6] The original algorithm sets a constant threshold $\theta$ such that $\operatorname{\mathbf{Pr}}[x_{\textsc{Opt}}\geq\theta]=\frac{1}{\mathrm{e}}$, which corresponds to $F(x_{i})\approx 1-\frac{1}{n}$ for each $i$ in the i.i.d. case.
The single-threshold algorithm achieves a competitive ratio of $1-\frac{1}{\mathrm{e}}\approx 0.632$ for the prophet secretary problem in terms of MaxExp. Hence, it is $(1-\frac{1}{\mathrm{e}})$-consistent for MaxExp. However, it suffers from the same counterexample and is only 0-robust like Gilbert-Mosteller algorithm.
In sum, the threshold functions corresponding to existing algorithms fail to achieve nontrivial trade-offs between consistency and robustness. They either ignore the predicted prior and have pessimistic consistency (Dynkin’s algorithm) or rely on having an accurate prior without any robustness (Gilbert-Mosteller algorithm and single-threshold algorithm).
3.2 Robust Threshold Functions
We next explain our design of robust threshold functions. First, we need to understand how an adversary would choose a worst-case prior distribution $F$, given the predicted prior $\tilde{F}$, threshold function $\theta$, and the corresponding algorithm $\mathcal{A}\big(\tilde{F},\theta\big)$. The key observation is the following lemma.
Lemma 6 (From mispredicted prior to misspecified thresholds).
For any distributions $F$ and $\tilde{F}$, and any threshold function $\theta$, algorithms $\mathcal{A}\big(\tilde{F},\theta\big)$ and $\mathcal{A}(F,F\circ\tilde{F}^{-1}\circ\theta)$ are the same.
Proof.
The lemma holds because $\tilde{F}(x)\,{>}\,\theta(t)$ if and only if $F(x)\,{>}\,F\circ\tilde{F}^{-1}\circ\theta(t)$, according to the monotonicity of $F$ and $\tilde{F}^{-1}$ and their boundary conditions. ∎
Despite its simple one-sentence proof, the lemma has an important implication. Instead of considering algorithm $\mathcal{A}\big(\tilde{F},\theta\big)$ parameterized with a mispredicted prior $\tilde{F}$, we may equivalently consider algorithm $\mathcal{A}\big(F,\theta\big)$ parameterized with the correct prior $F$ and a misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta$.
Importantly, the misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta$ is not arbitrary. By the boundary conditions of $F$ and $\tilde{F}^{-1}$, if the original threshold $\theta(t)$ equals 0 or $1$ at time $t$, the misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta(t)$ must have the same value 0 or $1$. Further, it is easy to show that this is the only constraint. Let $\lambda_{1}=\sup\{0\leq t\leq 1:\theta(t)=1\}$ and $\lambda_{2}=\inf\{0\leq t\leq 1:\theta(t)=0\}$. Then, the misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta(t)$ can be any non-increasing function for $t$ between $\lambda_{1}$ and $\lambda_{2}$, while having value $1$ for $0\leq t<\lambda_{1}$ and value 0 for $\lambda_{2}<t\leq 1$.
We will further show in Section 5 that the worst-case threshold $F\circ\tilde{F}^{-1}\circ\theta(t)$ is either always 0 or always $1$ in $\lambda_{1},\lambda_{2}]$. They correspond to a variant of Dynkin’s algorithm that starts to accept best-so-far values at a suboptimal time $\lambda=\lambda_{1}$ or $\lambda_{2}$, rather than the optimal choice $\frac{1}{\mathrm{e}}$. The competitive ratio $\lambda\ln\lambda$ of this algorithm (e.g., feldmanImprovedCompetitiveRatios2011) motivates the following definition of robust threshold functions.
Definition 5.
Consider any feasible robustness parameter $0\leq\beta\leq\frac{1}{\mathrm{e}}$. Let $0\leq\lambda_{1}\leq\frac{1}{\mathrm{e}}\leq\lambda_{2}\leq 1$ be the solutions to:
$\lambda\ln\lambda~=~\beta~.$
Readers familiar with the Lambert $\mathcal{W}$ function may note that:
$\lambda_{1}\coloneqq\mathrm{e}^{\mathcal{W}_{-1}(-\beta)}~,\quad\qquad\lambda_{2}\coloneqq\mathrm{e}^{\mathcal{W}_{0}(-\beta)}~.$ (1)
For any threshold function $\theta$, define a corresponding robust threshold function as:
$\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}(t)\coloneqq\begin{cases}1&0\leq t\leq\lambda_{1}~;\\ \theta(t)&\lambda_{1}<t<\lambda_{2}~;\\ 0&\lambda_{2}\leq t\leq 1~.\end{cases}$
For example, Figure 2 shows the robust counterpart of the threshold function from Gilbert-Mosteller algorithm.
Figure 2: Illustration of the robust threshold function corresponding to Gilbert–Mosteller algorithm, with $n=10$, $\beta=\nicefrac{{1}}{{3}}$, $\lambda_{1}=\mathrm{e}^{\mathrm{e}^{\mathcal{W}_{-1}(-\beta)}(-\nicefrac{{1}}{{3}})}\approx 0.220$ and $\lambda_{2}=\mathrm{e}^{\mathrm{e}^{\mathcal{W}_{0}(-\beta)}(-\nicefrac{{1}}{{3}})}\approx 0.538$.
3.3 Brief Discussion on Algorithms Outside the Framework
In terms of MaxExp, the literature has considered algorithms that accept the first value greater than a threshold at its arrival time, regardless of whether the value is best-so-far. With a correct prior, the optimal policy by backward induction takes this form, where the threshold is the expected accepted value if the algorithm rejects the current value. For example, see the work by krengel1977semiamarts on general prophet inequality and the more recent research [correaPostedPriceMechanisms2017, liuVariableDecompositionProphet2021, harbNewProphetInequalities2025] for i.i.d. prophet inequality. These algorithms are outside our framework, where the taken value must be best-so-far, in addition to surpassing the time-dependent threshold.
We stress that the best-so-far requirement is critical for robustness. The mentioned algorithms that have no best-so-far requirements are only 0-robust. To see this, consider a predicted prior supported on values much smaller than those in the real prior. Then, all values from the real prior will have cdf value essentially $1$ w.r.t. the predicted prior. As a result, the algorithm accepts the first value arrived when $\theta(t)<1$, which is noncompetitive and leads to 0 robustness.
Meanwhile, our framework is limited in that it does not adjust thresholds adaptively based on the observed values. Intuitively, a “learning” algorithm may adjust thresholds toward Dynkin’s algorithm [dynkinOptimumChoiceInstant1963] if the observed values are unlikely to be from the predicted prior, and conversely, toward the $0.5801$-competitive algorithm [gilbertRecognizingMaximumSequence] in the MaxProb setting or the $0.745$-competitive algorithm [correaProphetInequalitiesIID2019] in the MaxExp setting if the observed values are consistent with the predicted prior. We consider a static threshold function to make the analyses tractable. Exploring algorithms with adaptive thresholds is an interesting direction for future work.
4 Implicit Sharding
This section introduces an implicit sharding technique that reduces instances with a smaller number of values to instances with arbitrarily many values. In this sense, it suffices to design algorithms for the limit case when the number of values $n$ tends to infinity. The main result of this section is the following theorem.
Theorem 7.
For any $n,k\in\mathbb{N}$, and for both the MaxExp and MaxProb objectives, if there is a threshold function $\theta$ such that algorithm $\mathcal{A}\big(\tilde{F},\theta\big)$ is $\alpha$-consistent and $\beta$-robust for $nk$ values, then there is a corresponding $\alpha$-consistent $\beta$-robust algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ for $n$ values.
By our convention, $\tilde{F}^{\nicefrac{{1}}{{k}}}$ denotes the distribution defined by cdf $\tilde{F}^{\nicefrac{{1}}{{k}}}(\cdot)$. Importantly, the maximum of $k$ values drawn i.i.d. from $\tilde{F}^{\nicefrac{{1}}{{k}}}$ follows distribution $\tilde{F}$. We will construct an algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ for $n$ values to mimic the decision of algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ with $nk$ values drawn i.i.d. from $\tilde{F}^{\nicefrac{{1}}{{k}}}$. Let $x_{1},\dots,x_{n}$ be the $n$ values by their arrival order in the original instance. Further, consider an instance of $nk$ values $\{x^{\prime}_{ij}\}_{i\in[n],j\in[k]}$ drawn i.i.d. from distribution $F^{\nicefrac{{1}}{{k}}}$ with sorted arrival times $t^{\prime}_{11}<\cdots<t^{\prime}_{1k}<\cdots<t^{\prime}_{n1}<\cdots<t^{\prime}_{nk}$ drawn i.i.d. and uniformly from $[0,1]$.
Definition 6 (Coupling for Sharding).
Let $x_{i}=\max_{j\in[k]}x^{\prime}_{ij}$, and let $s_{i}$ be the arrival time of $\operatorname*{arg\,max}_{j\in[k]}x^{\prime}_{ij}$. We use $s_{i}$ instead of $t_{i}$ to avoid confusion, as $s_{i}$ is not uniform over $[0,1]$.
Naïvely, we would like to simulate algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ by sampling $x^{\prime}_{ij}$, $j\in[k]$, from $F^{\nicefrac{{1}}{{k}}}$ conditioned on the realization of $x_{i}$, and accept value $x_{i}$ if and only if one of the $x^{\prime}_{ij}$’s is accepted by $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$. This is the sharding technique by harbNewProphetInequalities2025 for the prophet inequality setting. Unfortunately, the algorithm does not know $F$ in the problem of optimal stopping with a predicted prior, and thus cannot sample $x^{\prime}_{ij}$ from $F^{\nicefrac{{1}}{{k}}}$. Sampling $x^{\prime}_{ij}$ based on the predicted prior $\tilde{F}$ and the corresponding $\tilde{F}^{\nicefrac{{1}}{{k}}}$ preserves the consistency of $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$, but not robustness.
We propose an implicit sharding method that partially simulates algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ only at the arrival times of the $x_{i}=\max_{j\in[k]}x^{\prime}_{ij}$. We will show that this partial simulation is at least as good as the full simulation in the original sharding method. Further, the partial simulation does not rely on $F$ and will not explicitly construct the $x^{\prime}_{ij}$’s, as they are only used in the analysis.
Bi-criteria Algorithm with Implicit Sharding $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ 1. Sample nknk arrival times i.i.d. from $\mathrm{Unif}[0,1]$ and sort them as:
             $t^{\prime}_{11}<\cdots<t^{\prime}_{1k}<\cdots<t^{\prime}_{n1}<\cdots<t^{\prime}_{nk}$ .
2. For each $i\in[n]$, sample $s_{i}$ uniformly from $\{t^{\prime}_{i1},t^{\prime}_{i2},\ldots,t^{\prime}_{ik}\}$. 3. For each value $x_{i}$ when no value has been accepted, accept $x_{i}$ if: (a) Value $x_{i}$ is best-so-far; and (b) The cdf value of $x_{i}$ w.r.t. $\tilde{F}^{\nicefrac{{1}}{{k}}}$ is greater than the threshold at time $s_{i}$, i.e., if:
              $\tilde{F}^{\nicefrac{{1}}{{k}}}(x_{i})>\theta(s_{i})~.$
Proof of Theorem 7.
It suffices to show that for any real prior $F$, and any realization of the values and arrival times under the defined coupling, the value accepted by algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ is greater than or equal to the value accepted by algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$.
Suppose algorithm $\mathcal{A}_{\mathrm{sharding}}(\tilde{F},\theta)$ accepts value $x_{i}$, and $x^{\prime}_{ij}=x_{i}$ is the corresponding value in the coupled instance for algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$. Since $x_{i}$ is best-so-far, $x^{\prime}_{ij}$ is also best-so-far because
$x^{\prime}_{ij}=x_{i}\geq x^{\prime}_{i^{\prime}j^{\prime}}\quad\text{for any }i^{\prime}\leq i,\;j^{\prime}\in[k].$
Further, the cdf value of $x^{\prime}_{ij}$ is greater than the threshold at time $t^{\prime}_{ij}$, as
$\tilde{F}^{\nicefrac{{1}}{{k}}}(x^{\prime}_{ij})=\tilde{F}^{\nicefrac{{1}}{{k}}}(x_{i})>\theta(s_{i})=\theta(t^{\prime}_{ij})~.$
Hence, algorithm $\mathcal{A}(\tilde{F}^{\nicefrac{{1}}{{k}}},\theta)$ accepts either $x^{\prime}_{ij}=x_{i}$, or a value that arrives earlier, which must be at most $x^{\prime}_{ij}=x_{i}$ since $x_{i}$ is best-so-far. Taking expectations over the coupling completes the proof. ∎
5 Proof of Robustness
We analyze the robustness of our algorithm for MaxProb (which then implies robustness for MaxExp) in a stronger setting that fixes the realized values and randomizes only the arrivals. This problem is also known as the game of Googol [fergusonWhoSolvedSecretary1989, gnedin1994solution].
Proposition 8.
For any predicted prior $\tilde{F}$, any threshold function $\theta$, any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, and any realized values $x_{1},x_{2},\dots,x_{n}$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ accepts the maximum value with probability at least
$\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}~,$
over the random arrivals of the values.
This implies the robustness guarantees for both MaxProb and MaxExp as corollaries through a standard argument.
Corollary 9.
For any predicted prior $\tilde{F}$, any threshold function $\theta$, and any $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ is $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$-robust for both MaxProb and MaxExp.
Proof.
By the law of total probability, we have
$\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]~=~\operatorname{\mathbf{E}}_{x_{1},\dots,x_{n}}\Big[\,\operatorname{\mathbf{Pr}}_{t_{1},\dots,t_{n}}\big[\,x_{\textsc{Alg}}=x_{\textsc{Opt}}\mid x_{1},\dots,x_{n}\,\big]\,\Big]~.$
By Proposition 8, the inner probability is at least $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$ for every realization, so the expectation is at least the same quantity. This proves robustness for MaxProb. Moreover, it implies $\operatorname{\mathbf{E}}[x_{\textsc{Alg}}]\geq\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}\cdot\operatorname{\mathbf{E}}[x_{\textsc{Opt}}]$, i.e., the same robustness for MaxExp. ∎
Therefore, it remains to prove Proposition 8. We first derive an explicit expression for the probability that algorithm $\mathcal{A}(\tilde{F},\theta)$ accepts the maximum value. Without loss of generality (by relabeling the distinct realized values), assume $x_{1}<x_{2}<\dots<x_{n}$, so $x_{\textsc{Opt}}=x_{n}$.
Lemma 10.
For any $n\in\mathbb{N}$, any predicted prior $\tilde{F}$, any threshold function $\theta$, any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, and any realized values $x_{1}<x_{2}<\dots<x_{n}$, algorithm $\mathcal{A}\big(\tilde{F},\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ accepts the maximum value $x_{n}$ with probability
$\int_{0}^{1}\mathds{1}{\scriptstyle{\{\tilde{F}(x_{n})>\theta(t)\}}}\bigg((1-t)^{n-1}+\sum_{i=1}^{n-1}(1-t)^{n-1-i}\int_{0}^{t}\mathds{1}{\scriptstyle{\{\tilde{F}(x_{i})\leq\theta(s)\}}}\,\mathrm{d}s\bigg)\,\mathrm{d}t~.$ (2)
Proof.
Since $x_{n}$ is always best-so-far, the algorithm accepts $x_{n}$ if and only if 1) its cdf value under $\tilde{F}$ is greater than the threshold, i.e., $\tilde{F}(x_{n})>\theta(t_{n})$, and 2) the algorithm accepts no value before $t_{n}$. By Proposition 4, the latter is equivalent to the event that the prefix maximum for time $t_{n}$ has cdf under $\tilde{F}$ at most the threshold at its arrival time. Hence, the acceptance probability is
$\int_{0}^{1}\mathds{1}{\scriptstyle{\{\tilde{F}(x_{n})>\theta(t_{n})\}}}\operatorname{\mathbf{Pr}}\big[\,\text{prefix maximum for time $t_{n}$ is below threshold}\,\big]\,\mathrm{d}t_{n}~.$
To compute the expression for the probability in the integrand, consider each possible prefix maximum separately. First, the prefix maximum is 0 if no value arrives before $t_{n}$, which occurs with probability $(1-t_{n})^{n-1}$. Further, the prefix maximum is $x_{i}$ if $x_{i}$ arrives at time $t_{i}\in[0,t_{n})$, and the larger values $x_{i+1},\ldots,x_{n-1}$ arrive after time $t_{n}$, which happens with probability $(1-t_{n})^{n-1-i}$. In this case, we also require $\tilde{F}(x_{i})\leq\theta(t_{i})$ for the arrival time $t_{i}\in[0,t_{n})$, contributing
$(1-t_{n})^{n-1-i}\int_{0}^{t_{n}}\mathds{1}{\scriptstyle{\{\tilde{F}(x_{i})\leq\theta(t_{i})\}}}\,\mathrm{d}t_{i}~.$
Putting together and changing variables $t_{n}$ with $t$ and $t_{i}$ with $s$, the lemma follows. ∎
We are now ready to prove Proposition 8.
Proof of Proposition 8.
By Lemma 10, it suffices to show that the expression in Equation 2 is at least $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$ for any predicted prior $\tilde{F}$ and threshold function $\theta$.
First, we fix $\theta$ and characterize the worst-case predicted prior $\tilde{F}$. Equation 2 depends on $\tilde{F}$ through $n$ values $0\leq\tilde{F}(x_{1})\leq\dots\leq\tilde{F}(x_{n})\leq 1$. Further, it is non-decreasing in $\tilde{F}(x_{n})$, and non-increasing in $\tilde{F}(x_{1}),\dots,\tilde{F}(x_{n-1})$. Therefore, given any fixed threshold function $\theta$, Equation 2 is minimized when all $\tilde{F}(x_{i})$ are equal to some $q\in[0,1]$.
In this case, the threshold function is effectively binary. A threshold $\theta(t)<q$ is equivalent to $\theta(t)=0$, as $\tilde{F}(x_{1}),\dots,\tilde{F}(x_{n})$ all surpass the threshold. Similarly, a threshold $\theta(t)\geq q$ is equivalent to $\theta(t)=1$, as none surpasses it. Hence, we may without loss of generality consider a binary threshold, such that for some $\lambda\in[\lambda_{1},\lambda_{2}]$, it holds
$\theta(t)=\begin{cases}1,&\text{if }t\leq\lambda~;\\ 0,&\text{otherwise~.}\end{cases}$
This corresponds to Dynkin’s algorithm with a suboptimal transition point of $\lambda$ instead of the optimal $\frac{1}{\mathrm{e}}$. This is $(-\lambda\ln\lambda)$-competitive (e.g., feldmanImprovedCompetitiveRatios2011). Since $\lambda\ln\lambda$ is concave in $\lambda$ on $[0,1]$, its minimum over $\lambda_{1},\lambda_{2}]$ is attained at an endpoint, yielding the bound $\min\{-\lambda_{1}\ln\lambda_{1},-\lambda_{2}\ln\lambda_{2}\}$. ∎
6 Algorithm for MaxExp: Proof of Theorem 1
By the implicit sharding method (Theorem 7), it suffices to design an algorithm $\mathcal{A}(\tilde{F},\theta)$ under the bi-criteria framework, and prove the desired consistency and robustness guarantees as the number of values $n$ tends to infinity. Further, by Corollary 9, the robustness guarantee holds for any robust threshold $\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ where $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$.
The rest of the section focuses on the consistency guarantee, for any fixed robustness parameter $\beta\in[0,\frac{1}{e}]$ and the corresponding $\lambda_{1},\lambda_{2}$. Since the predicted prior is correct, i.e., $\tilde{F}=F$, we will only use $F$ throughout the section. To formally study the limit as $n\to\infty$, we will design the threshold function for the extremal distribution $F^{n}$. More precisely, we will design a threshold function $\theta$ and analyze the consistency of algorithm $\mathcal{A}(F,\theta^{\nicefrac{{1}}{{n}}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ for $n$ values, and in particular, the limit when $n$ tends to infinity.
As a standard approach, we design the algorithm so that for any $\ell\geq 0$, it holds
$\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Alg}}\geq\ell\,\big]~\geq~\alpha\cdot\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Opt}}\geq\ell\,\big]~=~\alpha\cdot\big(1-F(\ell)^{n}\big)~.$ (3)
This implies the consistency guarantee as
$\operatorname{\mathbf{E}}\big[\,x_{\textsc{Alg}}\,\big]~=~\int_{0}^{\infty}\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Alg}}\geq\ell\,\big]\,\mathrm{d}\ell~\geq~\alpha\cdot\int_{0}^{\infty}\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Opt}}\geq\ell\,\big]\,\mathrm{d}\ell~=~\alpha\cdot\operatorname{\mathbf{E}}\big[\,x_{\textsc{Opt}}\,\big]~.$
6.1 Sufficient Conditions for Infinitely Many Values
We consider a general threshold function $\theta$ and derive an expression for the left-hand side of Equation 3, i.e., the probability that the algorithm accepts a value that is at least $\ell$.
Lemma 11.
For any threshold function $\theta$, any number of values $n$, and any $\ell\geq 0$, algorithm $\mathcal{A}\big(F,\theta^{\nicefrac{{1}}{{n}}}\big)$ satisfies that
$\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Alg}}\geq\ell\,\big]~=~\int_{F(\ell)^{n}}^{1}\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\Big(1-t+t\cdot\min\big\{\theta(s),q\big\}^{\nicefrac{{1}}{{n}}}\Big)^{n-1}q^{-\frac{n-1}{n}}\,\mathrm{d}s\,\mathrm{d}t\,\mathrm{d}q~,$
where $\theta^{-1}(x)\coloneqq\inf\{t\in[0,1]\colon\theta(t)<x\}$ is the generalized inverse of $\theta$, defined as the boundary above which $\theta(\cdot)$ is strictly smaller than $x$.
Proof.
Consider any value $x_{i}$ arriving at time $t_{i}$, with the prefix maximum for time $t_{i}$ being $y$ and arriving at time $s$. By Corollary 5, the algorithm $\mathcal{A}\big(F,\theta^{\nicefrac{{1}}{{n}}}\big)$ accepts $x_{i}$ if and only if
$F(x_{i})>\theta(t_{i})^{\nicefrac{{1}}{{n}}}~,\quad x_{i}\geq y~,\quad F(y)\leq\theta(s)^{\nicefrac{{1}}{{n}}}~.$
We consider the probability that $x_{i}$ is at least $\ell$ and accepted by the algorithm, enumerating over all triples $(x_{i},t_{i},s)$. First, value $x_{i}$ follows distribution $F$ and satisfies $x_{i}\geq\ell$. Further, its arrival time $t_{i}$ is uniform over $[0,1]$, and satisfies $F(x_{i})>\theta(t_{i})^{\nicefrac{{1}}{{n}}}$, or equivalently, $F(x_{i})^{n}>\theta(t_{i})$. Finally, given $x_{i}$ and $t_{i}$, the arrival time $s$ of the prefix maximum is uniformly distributed over $0,t_{i}]$. We may artificially sample $s$ from this distribution even if no values arrive before $t_{i}$. In sum, value $x_{i}$ is at least $\ell$ and accepted by the algorithm with probability
$\int_{x_{i}=\ell}^{x_{i}=\infty}\int_{\theta^{-1}(F(x_{i})^{n})}^{1}\int_{0}^{t_{i}}\frac{1}{t_{i}}\,\operatorname{\mathbf{Pr}}\big[\,y\leq x_{i}\,\,,\,F(y)\leq\theta(s)^{\nicefrac{{1}}{{n}}}\mid x_{i},t_{i},s\,\big]\,\mathrm{d}s\,\mathrm{d}t_{i}\,\mathrm{d}F(x_{i})~.$
For each of the n1n-1 values other than $x_{i}$, either it arrives after time $t_{i}$, in which case there is no constraint on the realized value, or it arrives before $t_{i}$, in which case the cdf of the realized value is at most $\min\{\theta(s)^{\nicefrac{{1}}{{n}}},F(x_{i})\}$ by the upper bounds for $y$. Thus,
$\operatorname{\mathbf{Pr}}\big[\,y\leq x_{i}\,\,,\,F(y)\leq\theta(s)^{\nicefrac{{1}}{{n}}}\mid x_{i},t_{i},s\,\big]~=~\big(\,1-t_{i}+t_{i}\min\{\theta(s)^{\nicefrac{{1}}{{n}}},F(x_{i})\}\,\big)^{n-1}~.$
Putting together and changing variables with $t=t_{i}$ and $q=F(x_{i})^{n}$, the above probability is
$\frac{1}{n}\int_{F(\ell)^{n}}^{1}\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\Big(1-t+t\cdot\min\big\{\theta(s),q\big\}^{\nicefrac{{1}}{{n}}}\Big)^{n-1}q^{-\frac{n-1}{n}}\,\mathrm{d}s\,\mathrm{d}t\,\mathrm{d}q~.$
Multiplying by $n$ by symmetry across all $x_{i}$’s proves the lemma. ∎
For any $y\in[0,1]$, if we let $n$ tend to infinity while keeping the same $F(\ell)^{n}=y$, the desired Inequality (3) becomes
$\int_{y}^{1}\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\theta(s),q\big\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t\,\mathrm{d}q~\geq~\alpha\cdot(1-y)~,$
where we use the fact that $\big(1-t+tz^{\nicefrac{{1}}{{n}}}\big)^{n-1}\to z^{t}$ as $n$ approaches infinity. [^7] Rigorously, suppose we have $\alpha^{*}$-consistency in the limit when $n$ tends to infinity and consider any $\varepsilon>0$. By the uniform convergence of the stated limit for $q\in[\varepsilon,1]$, we can apply implicit sharding to ensure Equation 3 for $\alpha=\alpha^{*}(1-\varepsilon)$ and any $q=F(\ell)^{n}\in[\varepsilon,1]$. Finally, Equation 3 holds for $\alpha^{*}(1-\varepsilon)^{2}$ when $q=F(\ell)^{n}\in[0,\varepsilon]$ because the left-hand-side is decreasing in $q$, while the right-hand-side differs by at most a $1-\varepsilon$ factor.
Define the function with support set $[0,1]$ as
$g(q)~\coloneqq~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\theta(s),q\big\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t~.$
Then, it suffices to design a threshold function $\theta$ such that for any $q\in[0,1]$, it holds
$g(q)~\geq~\alpha~.$
6.2 Proof of Theorem 1
Next, we focus on a robust threshold function $\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$, and for ease of notation we continue to write $\theta$ instead of $\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$, restricting $\theta$ from now on such that
  1. 1)
    $\theta(t)=1$ for $t\in[0,\lambda_{1})$; and
  2. 2)
    $\theta(t)=0$ for $t\in(\lambda_{2},1]$.
Further, recall the assumed properties in Theorem 1:
  1. 3)
    $\theta$ is left-continuous, which implies $\theta(\lambda_{1})=1$, and non-increasing; and
  2. 4)
    For any $z\in[\lambda_{1},\lambda_{2}]$, we have
    $\int_{z}^{1}\int_{0}^{t}\frac{1}{t}\cdot\theta(\max\{s,z\})^{t}\,\mathrm{d}s\,\mathrm{d}t~\geq~\alpha\cdot\theta(z)~.$
Consider any $q$ with $\theta(\lambda_{2})\leq q\leq\theta(\lambda_{1})=1$. Let $z=\theta^{-1}(q)$, and then by property 4), we have
$\alpha$ $~\leq~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\theta(\max\{s,\theta^{-1}(q)\})^{t}}{\theta(\theta^{-1}(q))}\,\mathrm{d}s\,\mathrm{d}t$
$~=~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\theta(s),\theta(\theta^{-1}(q))\big\}^{t}}{\theta(\theta^{-1}(q))}\,\mathrm{d}s\,\mathrm{d}t.$
By left-continuity of $\theta$ and the definition of $\theta^{-1}$, we have $\theta(\theta^{-1}(q))\geq q$. Therefore,
$\alpha~\leq~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\{\theta(s),q\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t~=~g(q).$
Finally, we verify $g(q)\geq\alpha$ for $q\in[0,\theta(\lambda_{2}))$ by reducing it to the first case.
Lemma 12.
For any $q\in[0,\theta(\lambda_{2})]$, we have
$g(q)\geq g\big(\theta(\lambda_{2})\big).$
Proof.
For any $q$ in the range $0,\theta(\lambda_{2})]$, we have $\theta^{-1}(q)=\lambda_{2}$. Also, observe that $\theta(s)\geq\theta(\lambda_{2})\geq q$ for $s\in[0,\lambda_{2}]$, and $\theta(s)=0$ for $s\in[\lambda_{2},1]$. Hence,
$g(q)=\int_{\lambda_{2}}^{1}\int_{0}^{\lambda_{2}}\frac{1}{t}q^{-(1-t)}\,\mathrm{d}s\,\mathrm{d}t=\int_{\lambda_{2}}^{1}\frac{\lambda_{2}}{t}q^{-(1-t)}\,\mathrm{d}t~,$
which is decreasing in $q$. ∎
7 Algorithm for MaxProb: Proof of Theorem 2
We first derive an explicit expression for the probability that algorithm $\mathcal{A}(\tilde{F},\theta)$ accepts the maximum value when the predicted prior is correct.
Lemma 13.
For any $n\in\mathbb{N}^{+}$, any prior $F$, and any threshold function $\theta$, algorithm $\mathcal{A}(\tilde{F}=F,\theta)$ is
$\Gamma_{n}(\theta)\coloneq\int_{0}^{1}\bigg(\int_{s}^{1}\frac{(1-t+t\theta(s))^{n}-t\theta(s)^{n}}{t(1-t)}\,\mathrm{d}t\;-\;\theta(s)^{n}\bigg)\,\mathrm{d}s$
competitive for MaxProb.
Proof.
Consider any value $x_{i}$ and the probability that $x_{i}$ is both maximum and accepted by the algorithm. Recall that $t_{i}$ is the arrival time of $x_{i}$, and let $s$ denote the arrival time of the prefix maximum $y$ for time $t_{i}$. By Corollary 5, the algorithm accepts $x_{i}$ if and only if
$F(x_{i})>\theta(t_{i})~,\quad x_{i}\geq y~,~\text{and}\quad F(y)\leq\theta(s)~.$
We enumerate over all possible triples $(x_{i},t_{i},s)$. First, arrival time $t_{i}$ distributes uniformly over $[0,1]$. Further, value $x_{i}$ follows distribution $F$, and must satisfy $F(x_{i})\geq\theta(t_{i})$ for $x_{i}$ to be accepted. Finally, given $x_{i}$ and $t_{i}$, the arrival time $s$ of the prefix maximum is uniformly distributed over $[0,t]$. We may artificially sample $s$ from this distribution even if no values arrive before $t_{i}$. Therefore, the probability that $x_{i}$ is both maximum and accepted equals
$\int_{0}^{1}\int_{\theta(t_{i})}^{1}\int_{0}^{t_{i}}\frac{1}{t_{i}}\,\operatorname{\mathbf{Pr}}\big[\,y\leq x_{i}\,\,,\,F(y)\leq\theta(s)\,\,,\,x_{\textsc{Opt}}=x_{i}\mid x_{i},t_{i},s\,\big]\,\mathrm{d}s\,\mathrm{d}F(x_{i})\,\mathrm{d}t_{i}~.$
For each of the n1n-1 values other than $x_{i}$, either it arrives before time $t_{i}$, in which case its cdf value must be at most $\min\{\theta(s),F(x_{i})\}$ by the upper bounds for the prefix maximum $y$; or it arrives after $t_{i}$, in which case it must be at most $x_{i}$ since $x_{i}$ is maximum. Therefore,
$\operatorname{\mathbf{Pr}}\big[\,y\leq x_{i}\,\,,\,F(y)\leq\theta(s)\,\,,\,x_{\textsc{Opt}}=x_{i}\mid x_{i},t_{i},s\,\big]~=~\Big(\,t_{i}\min\big\{\theta(s),F(x_{i})\big\}+(1-t_{i})F(x_{i})\,\Big)^{n-1}~.$
Putting together, changing variables with $t=t_{i}$ and $q=F(x_{i})$, and multiplying by $n$ by symmetry across all $x_{i}$’s, the algorithm accepts the maximum value with probability
$n\int_{0}^{1}\int_{\theta(t)}^{1}\int_{0}^{t}\frac{1}{t}\,\big(\,t\min\{\theta(s),q\}+(1-t)q\,\big)^{n-1}\,\mathrm{d}s\,\mathrm{d}q\,\mathrm{d}t~.$
Changing the order of integration and simplifying yields the stated expression for $\Gamma_{n}(\theta)$. ∎
Lemma 14.
For any $n\in\mathbb{N}^{+}$ and any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, define threshold function $\theta_{n}^{*}$ such that for any $s\in[0,1]$, $\theta_{n}^{*}(s)$ is the unique solution to
$\int_{s}^{1}\frac{\big(1-t+t\theta_{n}^{*}(s)\big)^{n-1}-\theta_{n}^{*}(s)^{n-1}}{1-t}\,\mathrm{d}t~=~\theta_{n}^{*}(s)^{n-1}~.$ (4)
We have
$\Gamma_{n}\big(\theta_{n}^{*}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~\geq~-\lambda_{1}\ln\lambda_{1}+\int_{\lambda_{1}}^{\lambda_{2}}\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~,$
where equality holds when $n$ tends to infinity.
Note that $\theta_{n}^{*}$ is exactly the threshold function used in the continuous-time Gilbert-Mosteller algorithm (Example 2). We defer the proofs of some lemmas to Appendix A.
Proof.
By the expression for $\Gamma_{n}(\theta)$ in Lemma 13, we derive the definition of $\theta_{n}^{*}$ from the first-order condition for $\theta(t)$, i.e.,
[…]
Further, $\theta_{n}^{*}(s)$ is decreasing in $s$ because Equation 4 can be rewritten as
$\int_{s}^{1}\frac{\big(\tfrac{1-t}{\theta_{n}^{*}(s)}+t\big)^{n-1}-1}{1-t}\,\mathrm{d}t~=~1~,$ (5)
except for the boundary condition $\theta_{n}^{*}(1)=0$.
In sum, $\theta_{n}^{*}$ is a threshold function satisfying the first-order conditions for $t\in[0,1]$. Therefore, $\theta_{n}^{*}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ is a threshold function satisfying the first-order conditions for $t\in[\lambda_{1},\lambda_{2}]$ on one hand, and the restrictions of being $1$ for $t\in[0,\lambda_{1})$ and 0 for $t\in(\lambda_{2},1]$ on the other hand.
We next write the competitive ratio $\Gamma_{n}\big(\theta_{n}^{*}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ as
$\int_{0}^{\lambda_{1}}\bigg(\int_{s}^{1}\frac{1}{t}\,\mathrm{d}t-1\bigg)\,\mathrm{d}s+\int_{\lambda_{1}}^{\lambda_{2}}\bigg(\int_{s}^{1}\frac{(1-t+t\theta_{n}^{*}(s))^{n}-t\theta_{n}^{*}(s)^{n}}{t(1-t)}\,\mathrm{d}t-\theta_{n}^{*}(s)^{n}\bigg)\,\mathrm{d}s+\int_{\lambda_{2}}^{1}\frac{(1-t)^{n-1}}{t}\,\mathrm{d}s~.$
We now focus on the limit case when $n$ tends to infinity. This is without loss of generality, either by the implicit sharding technique, or by the fact that the above ratio is non-increasing in $n$ (see Lemma 19 in Section A.1). In particular, the latter means that the stated consistency and robustness guarantees also hold directly for the bi-criteria algorithm without implicit sharding.
When $n\to\infty$, the first term equals $\lambda_{1}\ln\lambda_{1}$ by basic calculus, and the last term tends to 0. It remains to show that, as $n\to\infty$, the integrand of the second term converges to
$\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t~.$
By the definition of $\theta_{n}^{*}$ (Equation 4), the integrand equals
$\int_{s}^{1}\frac{(1-t+t\theta_{n}^{*}(s))^{n}-t\theta_{n}^{*}(s)^{n}}{t(1-t)}\,\mathrm{d}t\,-\,\theta_{n}^{*}(s)\int_{s}^{1}\frac{\big(1-t+t\theta_{n}^{*}(s)\big)^{n-1}-\theta_{n}^{*}(s)^{n-1}}{1-t}\,\mathrm{d}t$
$\qquad={}\int_{s}^{1}\frac{(1-t+t\theta_{n}^{*}(s))^{n-1}}{t}\,\mathrm{d}t~.$
Finally, we prove in Section A.2 (Lemma 20) that
$\theta_{n}^{*}(s)~=~\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o\!\left(\frac{1}{n}\right)$
uniformly on compact subsets of $[0,1)$, which yields the above limit as $n\to\infty$. [^8] Rigorously, for any $\varepsilon>0$, uniform convergence over $0,1-\varepsilon]$ implies the convergence of $\int_{s}^{1-\varepsilon}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t$. The difference from the ideal term is bounded by $\int_{1-\varepsilon}^{1}\frac{1}{t}\,\mathrm{d}t=\mathcal{O}(\varepsilon)$, which vanishes as $\varepsilon\to 0$.
8 Hardness for MaxProb
This section proves that certain consistency-robustness trade-offs for MaxProb are unattainable. We consider the discrete-time model and discrete distributions. In the presence of ties, we adopt the all-ties-win convention, i.e., the algorithm only needs to accept any value that equals $x_{\textsc{Opt}}$. This convention only strengthens the hardness result.
Theorem 15.
If there is an $\alpha$-consistent $\beta$-robust algorithm for MaxProb, then for any integers $n,K\in\mathbb{N}$ and any predicted prior $\tilde{F}$ over $[K]$, the following polytope $\mathcal{P}_{n,K,\tilde{F}}(\alpha,\beta)$ for variables $\{y_{t,\ell}\}_{t\in[n],\,\ell\in[K]}$ is feasible:
$\forall\ell,\quad$ $y_{0,\ell}=1;$ (6)
$\forall t,\forall\ell,\quad$ $0\leq y_{t,\ell}\leq 1;$ (7)
$\forall t,\forall\ell,\quad$ $\Delta^{t}(\ell)\,y_{t,\ell}-\Delta^{t-1}(\ell)\,\tilde{F}(\ell-1)\,y_{t-1,\ell}\geq 0$ (8)
$\forall t,\forall\ell,\quad$ $\Delta^{t}(\ell)\,y_{t,\ell}-\Delta^{t-1}(\ell)\,\tilde{F}(\ell-1)\,y_{t-1,\ell}\leq\sum_{m\in[\ell]}\Delta^{t-1}(m)\,\tilde{f}(\ell)\,y_{t-1,m};$ (9)
$\sum_{t\in[n]}\sum_{\ell\in[K]}\tilde{F}(\ell)^{n-t}\Bigg(\sum_{m\in[\ell]}\Delta^{t-1}(m)\,\tilde{f}(\ell)\,y_{t-1,m}+\Delta^{t-1}(\ell)\,\tilde{F}(\ell-1)\,y_{t-1,\ell}-\Delta^{t}(\ell)\,y_{t,\ell}\Bigg)\geq\alpha;$ (10)
$\forall k,\quad$ $\sum_{t\in[n]}\sum_{\ell\in[k]}\frac{\tilde{F}(\ell)^{n-t}}{\tilde{F}(k)^{n}}\Bigg(\sum_{m\in[\ell]}\Delta^{t-1}(m)\,\tilde{f}(\ell)\,y_{t-1,m}+\Delta^{t-1}(\ell)\,\tilde{F}(\ell-1)\,y_{t-1,\ell}-\Delta^{t}(\ell)\,y_{t,\ell}\Bigg)\geq\beta,$ (11)
where $\tilde{f}$ is the pmf of $\tilde{F}$, and we define $\Delta^{t}(\ell)\coloneqq\tilde{F}(\ell)^{t}-\tilde{F}(\ell-1)^{t}$ for any $t\in[n]$, with $\Delta^{0}(\ell)\coloneqq\mathds{1}{\scriptstyle{\{\ell=1\}}}$ under the convention $0^{0}=0$.
Additional Notations.
For any vector $\bm{x}=(x_{1},x_{2},\dots,x_{n})\in[K]^{n}$ and any $i,j\in[n]$, we define $\bm{x}_{i:j}\coloneqq(x_{i},x_{i+1}\dots,x_{j})$. Further define the maxima $\max\bm{x}_{i:j}=\max_{i\leq t\leq j}x_{t}$ and $\max\bm{x}=\max_{t\in[n]}x_{t}$. If $i>j$, we adopt the convention $\max\bm{x}_{i:j}=1$ consistent with the assumption that $x_{t}\geq 1$ for all $t$.
8.1 Family of Candidate Prior Distributions
Given the predicted prior $\tilde{F}$ over $[K]$, we consider a family of candidate prior distributions by conditioning on having values at most $k\in[K]$.
Definition 7.
Let $\mathcal{F}=\{F_{k}\}_{k\in[K]}$ be a family of distributions such that $F_{k}$ is the conditional distribution obtained by restricting $\tilde{F}$ to support $[k]$. In other words, the cdf of $F_{k}$ is
$F_{k}(\ell)~=~\begin{cases}\dfrac{\tilde{F}(\ell)}{\tilde{F}(k)},&1\leq\ell\leq k~;\\[8.61108pt] ~1,&k<\ell\leq K~.\end{cases}$
At the two extremes, we have $F_{K}=\tilde{F}$ and $F_{1}$ being a point mass at $1$. We use $\operatorname{\mathbf{Pr}}_{k}$ to denote probability under distribution $F_{k}$, and $\operatorname{\mathbf{Pr}}$ for probabilities of events independent of the specific choice of the underlying distribution, e.g., those over the randomness of the algorithm.
The main property of this construction is the following. Given any realized values in the first $t$ steps, the only information we can conclude about the real prior $F_{k}$ is $k\geq\max\bm{x}_{1:t}$. Formally, we have the next lemma.
Lemma 16.
For any $t\in[n]$, any $\ell\in[K]$, and any subset $S\subseteq[\ell]^{t}$ of $t$ realized values whose maximum equals $\ell$, the conditional probability:
$\operatorname{\mathbf{Pr}}_{k}\big[\bm{x}_{1:t}\in S\mid\max\bm{x}_{1:t}=\ell\big]$
is the same for all $k$ with $\ell\leq k\leq K$.
Proof.
By Bayes’ rule, we have
$\operatorname{\mathbf{Pr}}_{k}\big[\bm{x}_{1:t}\in S\mid\max\bm{x}_{1:t}=\ell\big]~=~\frac{\operatorname{\mathbf{Pr}}_{k}\big[\bm{x}_{1:t}\in S\,,\,\max\bm{x}_{1:t}=\ell\big]}{\operatorname{\mathbf{Pr}}_{k}\big[\max\bm{x}_{1:t}=\ell\big]}~.$
Replacing $k$ with $K$ corresponds to multiplying both the numerator and denominator by $\tilde{F}^{t}(k)$ by Definition 7, resulting in the same ratio. ∎
8.2 Representations of Algorithms
To derive the linear constraints in polytope $\mathcal{P}_{n,K,\tilde{F}}(\alpha,\beta)$, we consider two representations of online algorithms when the predicted prior is $\tilde{F}$ and the real prior is one of the conditional distributions in the above family $\mathcal{F}$. The first one relies on the conditional acceptance probabilities, which appeared in several previous works (e.g., buchbinderSecretaryProblemsLinear2014, correaSampledrivenOptimalStopping2021).
Definition 8 (Conditional Acceptance Probabilities).
Given any algorithm, and for any $t\in[n]$, define $\textsc{Acc}_{t}:[K]^{t}\to[0,1]$ as the probability of accepting the value at time $t$, conditioned on the realized values up to time $t$, and not accepting any value before $t$. That is
$\textsc{Acc}_{t}(\bm{x}_{1:t})~\coloneqq~\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}=t\mid\textsc{Alg}\geq t\,,\,\bm{x}_{1:t}=\bm{x}_{1:t}\,\big]~.$
The conditional acceptance probabilities fully characterize the algorithm, as we can write the probability of accepting the value at step $t$ as:
$\operatorname{\mathbf{Pr}}\big[\textsc{Alg}=t\mid\bm{x}\big]~=~\textsc{Acc}_{t}(\bm{x}_{1:t})\prod_{i\in[t-1]}\big(1-\textsc{Acc}_{i}(\bm{x}_{1:i})\big)~.$
Next, we show that we may without loss of generality focus on a class of minor-oblivious algorithms, whose decisions for whether accepting a best-so-far value $x_{t}$ at time $t$ is independent of smaller values $\bm{x}_{1:t-1}$ in the past.
Definition 9 (Minor-oblivious Algorithms).
An algorithm is called minor-oblivious if $\textsc{Acc}_{t}(\bm{x}_{1:t})=\textsc{Acc}_{t}(\bm{y}_{1:t})$ for any $\bm{x}_{1:t}$ and $\bm{y}_{1:t}$ satisfying $x_{t}=y_{t}=\max\bm{x}_{1:t}=\max\bm{y}_{1:t}$. We then write $\textsc{Acc}_{t}(\ell)$ for the common value of $\textsc{Acc}_{t}(\bm{x}_{1:t})$ for all $\bm{x}_{1:t}$ satisfying $\max\bm{x}_{1:t}=\ell$.
Recall that based on the realization of $\bm{x}_{1:t}$, the only information we may conclude about the prior $F_{k}$ is $k\geq\max\bm{x}_{1:t}$. Intuitively, there is no reason to make decisions based on the details of $\bm{x}_{1:t}$ other than a maximum $\max\bm{x}_{1:t}$. In this sense, minor-oblivious algorithms are without loss of generality. We formalize this characterization in the following proposition, which may be of independent interest. The proof is deferred to Section C.1.
Proposition 17.
If an algorithm is $\alpha$-competitive when the prior is $F_{K}$ and $\beta$-competitive when the prior is $F_{k}$ for $k\in[K-1]$, there is a minor-oblivious algorithm with the same guarantees.
Focusing on minor-oblivious algorithms, we next consider a new representation by the probability of rejecting the first $t$ values conditioned on a maximum among them.
Definition 10 (Minor-oblivious Rejection Probabilities).
For any $t\in[n]$ and any $\ell\in[K]$, define the minor-oblivious rejection probability $\textsc{Rej}_{t}:[K]\to[0,1]$ as:
$\textsc{Rej}_{t}(\ell)~\coloneqq~\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}>t\mid\max\bm{x}_{1:t}=\ell\,\big]~.$
We artificially define $\textsc{Rej}_{0}(\ell)\coloneqq\mathds{1}{\scriptstyle{\{\ell=1\}}}$ to simplify some expressions in the rest of the section.
There is a one-to-one correspondence between the two representations (see Section C.2). We state below the direction to be used in the proof of Theorem 15.
Lemma 18.
For any minor-oblivious rule, any $t\in[n]$, and any $\ell\in[K]$, we have
$\textsc{Acc}_{t}(\ell)~=~1-\frac{\textsc{Rej}_{t}(\ell)\cdot\operatorname{\mathbf{Pr}}_{K}\big[\max\bm{x}_{1:t}=\ell\,\big]-\textsc{Rej}_{t-1}(\ell)\cdot\operatorname{\mathbf{Pr}}_{K}\big[\,x_{t}<\max\bm{x}_{1:t-1}=\ell\,\big]}{\sum_{m\in[\ell]}\textsc{Rej}_{t-1}(m)\cdot\operatorname{\mathbf{Pr}}_{K}\big[\max\bm{x}_{1:t-1}=m,x_{t}=\ell\,\big]}~.$ (12)
8.3 Proof of Theorem 15
Given any minor-oblivious, $\alpha$-consistent, $\beta$-robust algorithm, we may assume without loss of generality that it rejects all values that are not best-so-far, since we consider MaxProb. We next verify that $y_{t,\ell}=\textsc{Rej}_{t}(\ell)$ is feasible w.r.t. polytope $\mathcal{P}_{n,K,\tilde{F}}(\alpha,\beta)$.
Bounds for Rejection Probabilities.
By the definition of the minor-oblivious rejection probabilities, and the artificial definition of $\textsc{Rej}_{0}(\ell)$, we obtain Constraints (6) and (7):
$\textsc{Rej}_{0}(\ell)\,=\,1~,\qquad 0\,\leq\,\textsc{Rej}_{t}(\ell)\,\leq\,1~,\qquad\forall t\in[n],\;\forall\ell\in[K]~.$
Bounds for Acceptance Probabilities.
By $0\leq\textsc{Acc}_{t}(\ell)\leq 1$ and Equation 12, we get that
$0$ $~\leq~\textsc{Rej}_{t}(\ell)\cdot\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t}=\ell]-\textsc{Rej}_{t-1}(\ell)\cdot\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t-1}=\ell,x_{t}<\ell]$
$~\leq~\sum_{m\in[\ell]}\textsc{Rej}_{t-1}(m)\cdot\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t-1}=m,x_{t}=\ell]~.$
These two inequalities correspond to Constraints (8) and (9), because
$\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t}=\ell]$ $~=~\Delta^{t}(\ell)~,$
$\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t-1}=\ell,x_{t}<\ell]$ $~=~\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t-1}=\ell]\cdot\operatorname{\mathbf{Pr}}_{K}[x_{t}<\ell]$
$~=~\Delta^{t-1}(\ell)\tilde{F}(\ell-1)~,$
$\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t-1}=m,x_{t}=\ell]$ $~=~\operatorname{\mathbf{Pr}}_{K}[\max\bm{x}_{1:t-1}=m]\cdot\operatorname{\mathbf{Pr}}_{K}[x_{t}=\ell]$
[…]
Competitive Ratios.
Consider the probability of accepting a maximum value when the prior is $F_{k}$. We write it as:
$\operatorname{\mathbf{Pr}}_{k}[x_{\textsc{Alg}}=x_{\textsc{Opt}}]$ $~=~\sum_{t\in[n]}\sum_{\ell\in[k]}\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}=t\,,\,\max\bm{x}_{1:t}=\max\bm{x}=\ell\,\big]$
$~=~\sum_{t\in[n]}\sum_{\ell\in[k]}\operatorname{\mathbf{Pr}}_{k}\big[\,\max\bm{x}_{1:t}=\max\bm{x}=\ell\,\big]\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}=t\mid\max\bm{x}_{1:t}=\ell\,\big]~.$
We did not explicitly write the optimality of $x_{t}$, i.e., $x_{t}=\max\bm{x}$, since $\textsc{Alg}=t$ implies $x_{t}=\max\bm{x}_{1:t}$ by our assumption without loss of generality that the algorithm rejects all values that are not best-so-far.
By the relation between $\tilde{F}=F_{K}$ and $F_{k}$, we can rewrite it as:
$\frac{1}{\tilde{F}(k)^{n}}\sum_{t\in[n]}\sum_{\ell\in[k]}\operatorname{\mathbf{Pr}}_{K}\big[\,\max\bm{x}_{1:t}=\max\bm{x}=\ell\,\big]\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}=t\mid\max\bm{x}_{1:t}=\ell\,\big]$
$\qquad\qquad~=~\frac{1}{\tilde{F}(k)^{n}}\sum_{t\in[n]}\sum_{\ell\in[k]}\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}=t\,,\,\max\bm{x}_{1:t}=\max\bm{x}=\ell\,\big]~.$
Further, the event $\textsc{Alg}=t$ only depends on $\bm{x}_{1:t}$, while the realization of $\bm{x}_{t+1:n}$ is independent to both $\bm{x}_{1:t}$ and whether $\textsc{Alg}=t$. We can further transform the above expression into:
$\frac{1}{\tilde{F}^{n}(k)}\sum_{t\in[n]}\sum_{\ell\in[k]}\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}=t\,,\,\max\bm{x}_{1:t}=\ell\,\big]\operatorname{\mathbf{Pr}}_{K}\big[\,\max\bm{x}_{t+1:n}\leq\ell\,\big]$
$\qquad\qquad~=~\sum_{t\in[n]}\sum_{\ell\in[k]}\frac{\tilde{F}(\ell)^{n-t}}{\tilde{F}(k)^{n}}\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}=t\,,\,\max\bm{x}_{1:t}=\ell\,\big]~.$ (13)
The final step is to write the above probability as a linear expression of the minor-oblivious rejection probabilities $\textsc{Rej}_{t}(\ell)$’s. First, we write $\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}=t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]$ as:
$\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}\geq t\,,\,\max\bm{x}_{1:t}=\ell\,\big]\,-\,\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}>t\,,\,\max\bm{x}_{1:t}=\ell\,\big]~.$ (14)
The first part equals:
$\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}\geq t\,,\,\max\bm{x}_{1:t-1}\leq\ell\,,\,x_{t}=\ell\,\big]+\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}\geq t\,,\,\max\bm{x}_{1:t-1}=\ell\,,\,x_{t}<\ell\,\big]$
$\qquad~=~\sum_{m\in[\ell]}\operatorname{\mathbf{Pr}}[\max\bm{x}_{1:t-1}=m]\cdot\operatorname{\mathbf{Pr}}[x_{t}=\ell]\cdot\operatorname{\mathbf{Pr}}\big[\textsc{Alg}>t-1\mid\max\bm{x}_{1:t-1}=m\big]$
$\qquad\qquad\qquad~+\operatorname{\mathbf{Pr}}[\max\bm{x}_{1:t-1}=\ell]\cdot\operatorname{\mathbf{Pr}}[x_{t}<\ell]\cdot\operatorname{\mathbf{Pr}}\big[\textsc{Alg}>t-1\mid\max\bm{x}_{1:t-1}=\ell\big]$
$\qquad~=~\sum_{m\in[\ell]}\Delta^{t-1}(m)\cdot\tilde{f}(\ell)\cdot\textsc{Rej}_{t-1}(m)+\Delta^{t-1}(\ell)\cdot\tilde{F}(\ell-1)\cdot\textsc{Rej}_{t-1}(\ell)~.$ (15)
The second part equals:
$\operatorname{\mathbf{Pr}}_{K}\big[\,\max\bm{x}_{1:t}=\ell\,\big]\cdot\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}>t\mid\max\bm{x}_{1:t}=\ell\,\big]~=~\Delta^{t}(\ell)\cdot\textsc{Rej}_{t}(\ell)~.$ (16)
Combining Equations 13, 14, 15 and 16, we conclude that when the real prior is $F_{k}$, the competitive ratio is equal to
$\sum_{t\in[n]}\sum_{\ell\in[k]}\frac{\tilde{F}(\ell)^{n-t}}{\tilde{F}(k)^{n}}\bigg($ $\sum_{m\in[\ell]}\Delta^{t-1}(m)\cdot\tilde{f}(\ell)\cdot\textsc{Rej}_{t-1}(m)$
$\quad+\Delta^{t-1}(\ell)\cdot\tilde{F}(\ell-1)\cdot\textsc{Rej}_{t-1}(\ell)-\Delta^{t}(\ell)\cdot\textsc{Rej}_{t}(\ell)\bigg).$
Since the algorithm is $\alpha$-consistent and $\beta$-robust, the above competitive ratio is at least $\alpha$ when $k=K$, and at least $\beta$ when $k\in[K-1]$. These correspond to Constraints (10) and (11).
8.4 Numerical Results
For any $\lambda\in[0,1]$, we solve a linear program with variables $\{y_{t,\ell}\}_{t\in[n],\,\ell\in[K]}\cup\{\alpha,\beta\}$ to maximize:
$\lambda\cdot\alpha+(1-\lambda)\cdot\beta$
subject to the linear constraints that define polytope $\mathcal{P}_{n,K,\,\tilde{F}}(\alpha,\beta)$. The optimal objective value $\mathrm{LP}^{*}(\lambda)$ then implies a hardness result of the form:
$\lambda\cdot\alpha+(1-\lambda)\cdot\beta~\geq~\mathrm{LP}^{*}(\lambda)~.$
We set $n=30$ and $K=1024$, and consider a predicted prior $\tilde{F}$ defined by $\tilde{f}(k)\propto\frac{1}{k}$ for $k\in[K]$. We solve the linear programs for $\lambda=\frac{i}{100}$ for $0\leq i\leq 100$ to get the curve shown in Figure 1(b). For details, see our code repository.
Appendix A Missing Proofs for MaxProb Consistency
A.1 Monotonicity of Consistency
Lemma 19.
For any fixed $\lambda_{1},\lambda_{2}\in[0,1]$, the consistency $\Gamma_{n}\big(\theta^{*}_{n}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ is non-increasing in $n$.
Proof.
We show that for any $n$, it holds:
$\Gamma_{n+1}\big(\theta^{*}_{n+1}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~\leq~\Gamma_{n}\big(\theta^{*}_{n}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~.$
Consider a relaxed version of the problem with $n+1$ values, where the adversary tells the algorithm at the beginning the information about the smallest realized value, including its arrival time $\tau$ and its value $x_{\tau}$. We consider two policies.
In the first policy, we ignore the additional information and directly apply the algorithm parameterized by $\theta^{*}_{n+1}$. By definition, the consistency of this policy is $\Gamma_{n+1}\big(\theta^{*}_{n+1}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$.
We now describe the second and optimal policy. Conditioned on $\tau$ and $x_{\tau}$, the remaining $n$ values are i.i.d. from $F$ restricted to $x_{\tau},+\infty)$, and their arrival times are i.i.d. uniform on $[0,1]$. Thus, by the optimality of $\theta^{*}_{n}$, any algorithm that rejects all values during $0,\lambda_{1}]$, accepts best-so-far values during $\lambda_{1},\lambda_{2}]$ based on a threshold function, and accepts best-so-far values during $\lambda_{2},1]$ has consistency at most $\Gamma_{n}\big(\theta^{*}_{n}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$.
Since the second strategy is optimal, we get the stated inequality. ∎
A.2 Asymptotic Characterization of $\theta_{n}^{*}$
Lemma 20.
For any fixed $s\in[0,1)$, we have
$\theta^{*}_{n}(s)~=~\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o\Big(\frac{1}{n}\Big)~,$
where $\gamma\approx 0.80435$ satisfies
$\sum_{k\geq 1}\frac{\gamma^{k}}{k!k}=1~.$
Proof.
Recall Equation 5:
$\int_{s}^{1}\frac{\big(\frac{1-t}{\theta^{*}_{n}(s)}+t\big)^{n-1}-1}{1-t}\,\mathrm{d}t=1~.$
Fix any $s$, substitute $x_{n}=\frac{1}{\theta^{*}_{n}(s)}-1$, the left-hand side becomes
$\int_{s}^{1}\frac{\left(1+x_{n}(1-t)\right)^{n-1}-1}{1-t}\,\mathrm{d}t~.$
Expanding the left-hand side with the binomial theorem, we have
$(n-1)(1-s)x_{n}+\binom{n-1}{2}\frac{(1-s)^{2}}{2}x_{n}^{2}+\cdots+\binom{n-1}{k}\frac{(1-s)^{k}}{k!}x_{n}^{k}+\cdots+\frac{(1-s)^{n-1}}{(n-1)!}x_{n}^{n-1}=1~.$
Now, we analyze the asymptotic behavior of $x_{n}$. Obviously, we have $x_{n}=\mathcal{O}(\tfrac{1}{n-1})$; otherwise, the left-hand side either diverges to infinity or converges to 0, thus cannot equal $1$.
Now, we find the constant such that $(n-1)(1-s)x_{n}\to c$ for some constant $c$ as $n\to\infty$. Note that such a constant $c$ shall satisfy
$c+\frac{n-2}{n-1}\cdot\frac{c^{2}}{2!2}+\cdots+\frac{(n-2)(n-3)\cdots(n-k)}{(n-1)^{k-1}}\cdot\frac{c^{k}}{k!k}+\cdots=1.$
Intuitively, as $n\to\infty$, the coefficients $\frac{(n-2)(n-3)\cdots(n-k)}{(n-1)^{k-1}}\to 1$ for any fixed $k$. Thus, we expect that $c$ satisfies
$\sum_{k\geq 1}\frac{c^{k}}{k!k}=1~,$
which defines the constant $\gamma\approx 0.80435$ in the lemma.
We can prove the convergence more rigorously by showing that, for any $c\in(0,1)$, the difference between the ideal and the actual sum tends to 0 as $n\to\infty$.
Consider the difference:
$\sum_{k=1}^{n}\frac{c^{k}}{k!k}\left(1-\left(1-\frac{1}{n-1}\right)\left(1-\frac{2}{n-1}\right)\cdots\left(1-\frac{k-1}{n-1}\right)\right)~.$
Observe that $1\leq(1-x_{1})(1-x_{2})\cdots(1-x_{m})+x_{1}+x_{2}+\cdots+x_{m}$ for any $x_{i}\in[0,1]$. Indeed, the partial derivative shows that the right-hand side is increasing in each $x_{i}$, thus the minimum is attained at $x_{i}=0$ for all $i$, which is $1$.
It follows that the difference is bounded by
$\sum_{k=1}^{n}\frac{c^{k}}{k!k}\cdot\frac{1+2+\cdots+(k-1)}{n-1}=\sum_{k=1}^{n}\frac{c^{k}}{k!k}\cdot\frac{k(k-1)}{2(n-1)}\leq\frac{1}{2(n-1)}\sum_{k=1}^{\infty}\frac{1}{(k-1)!}~.$
Using the fact that $\sum_{k=1}^{\infty}\frac{1}{(k-1)!}=\mathrm{e}$, the bound is $\mathcal{O}(\tfrac{1}{n})$ and therefore tends to 0 as $n\to\infty$.
Finally, we conclude that $(n-1)(1-s)x_{n}\to\gamma$ as $n\to\infty$, which implies that $x_{n}=\frac{\gamma}{(n-1)(1-s)}+o(\tfrac{1}{n})$. This completes the proof by $\theta^{*}_{n}(s)=\frac{1}{1+x_{n}}=\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o(\tfrac{1}{n})$. ∎
Appendix B Numerical Solution to the Differential Equation in Theorem 1
Recall the differential equation in Theorem 1:
$\theta(\lambda_{1})$ $~=~1~,$ (17)
$\int_{z}^{1}\int_{0}^{t}\frac{1}{t}\theta\big(\max\{s,z\}\big)^{t}\,\mathrm{d}s\,\mathrm{d}t$ $~\geq~\alpha\cdot\theta(z)~,\qquad\forall z\in[\lambda_{1},\lambda_{2}]~.$ (18)
We assume throughout that $\theta$ is left‑continuous, non‑increasing, and $\theta(s)\equiv 0$ for $s\in[\lambda_{2},1]$.
Define the left-hand side of Equation 18 as $L(z)$ and simplify:
$L(z)~:=$ $~\int_{z}^{1}\int_{0}^{t}\frac{1}{t}\theta\big(\max\{s,z\}\big)^{t}\,\mathrm{d}s\,\mathrm{d}t$ (19)
$~=$ $~\int_{z}^{1}\bigg(\frac{z\,\theta(z)^{t}}{t}+\int_{z}^{t}\frac{\theta(s)^{t}}{t}\,\mathrm{d}s\bigg)\,\mathrm{d}t$
$~=$ $~z\int_{z}^{1}\frac{\theta(z)^{t}}{t}\,\mathrm{d}t+\int_{z}^{\lambda_{2}}\int_{s}^{1}\frac{\theta(s)^{t}}{t}\,\mathrm{d}t\,\mathrm{d}s~.$
Here, the last equality is obtained by exchanging the order of integration and using $\theta\equiv 0$ on $\lambda_{2},1]$.
We approximate $\theta$ by a left‑continuous non‑increasing step function computed backward from $\lambda_{2}$ to $\lambda_{1}$ so that Equation 18 holds for all $z\in[\lambda_{1},\lambda_{2}]$.
Definition 11 (Step function approximation).
Given an integer $m>0$, define a uniform grid $\{z_{i}\}_{i=1}^{m+1}$ by $z_{i}=\lambda_{1}+\tfrac{i-1}{m}(\lambda_{2}-\lambda_{1})$ for $i=1,\dots,m+1$ (so $z_{m+1}=\lambda_{2}$). We say that $\theta$ is a step function with step values $\theta_{1},\dots,\theta_{m}$ if $\theta(z)=\theta_{i}$ for all $z\in(z_{i},z_{i+1}]$, and $\theta(\lambda_{1})=\theta_{1}$.
Lemma 21 (Monotonicity of $L(z)$).
$L(z)$ is decreasing in $z$ on $\lambda_{1},\lambda_{2}]$.
Proof.
As $z$ increases, the integration domain in Equation 19 shrinks with a non-negative integrand. The integrand is non-increasing since $\theta$ is non-increasing. ∎
Lemma 22 (Endpoint reduction).
A step function $\theta$ with $m$ steps satisfies Equation 18 for all $z\in[\lambda_{1},\lambda_{2}]$ if it satisfies Equation 18 at the endpoints of each interval, i.e., for all $i\in[m]$,
$L(z_{i+1})~\geq~\alpha\,\theta_{i}~.$
Proof.
By Lemma 21, $L(z)$ is decreasing in $z$. Thus, for any $z\in(z_{i},z_{i+1}]$, it holds $L(z)\geq L(z_{i+1})\geq\alpha\,\theta_{i}=\alpha\,\theta(z)$. ∎
Next, we describe how to compute the step values $\theta_{1},\dots,\theta_{m}$ given $\alpha$ and $m$ so that Lemma 22 holds.
Initialization at $z=\lambda_{2}$.
For $i=m$, Lemma 22 is
$\lambda_{2}\int_{\lambda_{2}}^{1}\frac{\theta(\lambda_{2})^{t}}{t}\,\,\mathrm{d}t~\geq~\alpha\,\theta(\lambda_{2})~.$
Letting $\theta_{m}\coloneqq\theta(\lambda_{2})\neq 0$, this is
$\lambda_{2}\int_{\lambda_{2}}^{1}\frac{\theta_{m}^{-(1-t)}}{t}\,\,\mathrm{d}t~\geq~\alpha~,$
where the left-hand side decreases from $\infty$ to 0 as $\theta_{m}$ increases from 0 to $\infty$. Thus, there exists a $\theta_{m}^{*}>0$ such that the equality holds. [^9] During experiment, we use binary search to locate $\theta_{m}^{*}$ with $100$ iterations so that the error can be ignored.
Backward recursion.
For $i=m-1,m-2,\ldots,1$, suppose we have computed $\theta_{i+1}^{*},\theta_{i+2}^{*},\ldots,\theta_{m}^{*}$. Using Equation 19, we have
$L(z_{i+1})~=$ $z_{i+1}\int_{z_{i+1}}^{1}\frac{\theta_{i}^{t}}{t}\,\,\mathrm{d}t+\sum_{j=i+1}^{m}\int_{z_{j}}^{z_{j+1}}\int_{s}^{1}\frac{{\theta^{*}_{j}}^{t}}{t}\,\,\mathrm{d}t\,\,\mathrm{d}s$
$~=$ $z_{i+1}\int_{z_{i+1}}^{1}\frac{\theta_{i}^{t}}{t}\,\,\mathrm{d}t+\sum_{j=i+1}^{m}\left[\int_{z_{j}}^{z_{j+1}}\frac{t-z_{j}}{t}\,{\theta_{j}^{*}}^{t}\,\,\mathrm{d}t+(z_{j+1}-z_{j})\int_{z_{j+1}}^{1}\frac{{\theta^{*}_{j}}^{t}}{t}\,\,\mathrm{d}t\right]~.$
Thus, Lemma 22 is simplified to
$z_{i+1}\int_{z_{i+1}}^{1}\frac{\theta_{i}^{t}}{t}\,\,\mathrm{d}t+\sum_{j=i+1}^{m}\left[\int_{z_{j}}^{z_{j+1}}\frac{t-z_{j}}{t}\,{\theta_{j}^{*}}^{t}\,\,\mathrm{d}t+(z_{j+1}-z_{j})\int_{z_{j+1}}^{1}\frac{{\theta^{*}_{j}}^{t}}{t}\,\,\mathrm{d}t\right]~\geq~\alpha\,\theta_{i}~.$ (20)
Equivalently,
$z_{i+1}\int_{z_{i+1}}^{1}\frac{\theta_{i}^{-(1-t)}}{t}\,\,\mathrm{d}t+\theta_{i}^{-1}\,\sum_{j=i+1}^{m}\left[\int_{z_{j}}^{z_{j+1}}\frac{t-z_{j}}{t}\,{\theta^{*}_{j}}^{t}\,\,\mathrm{d}t+(z_{j+1}-z_{j})\int_{z_{j+1}}^{1}\frac{{\theta^{*}_{j}}^{t}}{t}\,\,\mathrm{d}t\right]~\geq~\alpha~,$ (21)
whose left-hand side decreases from $\infty$ to 0 as $\theta_{i}$ increases from 0 to $\infty$. Again, there exists a $\theta_{i}^{*}>0$ such that the equality holds.
We select the largest feasible $\theta_{i}=\theta_{i}^{*}$ by binary search such that equality holds.
Lemma 23 (Monotonicity of the step values).
The sequence $(\theta_{i}^{*})_{i=1}^{m}$ produced by the recursion is non-increasing in $i$.
Proof.
By definition, $\alpha\theta_{i}^{*}=L(z_{i+1})\geq L(z_{i+2})=\alpha\theta_{i+1}^{*}$, where the inequality holds by Lemma 21. ∎
Enforcing $\theta(\lambda_{1})=1$.
We only need to find an $\alpha$ with the corresponding sequence $(\theta_{i}^{*})_{i=1}^{m}$ satisfying Equation 18 and $\theta_{1}^{*}\geq 1$. The following lemma shows this suffices.
Lemma 24.
If a step function $\theta$ with step values $\theta_{1},\theta_{2},\dots,\theta_{m}$ satisfies Equation 18, then the step function with step values $(\min\{\theta_{i},1\})_{i=1}^{m}$ also satisfies Equation 18.
Proof.
By unifying intervals that $\theta_{i}$ exceeds $1$, as Lemma 22, we only need to check that Equation 21 holds for $\min\{\theta_{i},1\}$, where $i$ is the last $\theta_{i}$ exceeding $1$. This instantly stands by the non-increasing of the left-hand side. ∎
A direct corollary implies the desired result.
Corollary 25.
If an $\alpha$ such that the corresponding recursively constructed sequence $(\theta_{i}^{*})_{i=1}^{m}$ satisfies $\theta_{1}^{*}\geq 1$, then the step function $\theta^{*}$ with step values $(\min\{\theta_{i}^{*},1\})_{i=1}^{m}$ is a left-continuous non-increasing function satisfying Equation 17 and Equation 18.
In our implementation, we fix several pairs of values $(\alpha,\beta)$ and numerically verify that for each pair, the constructed sequence $(\theta_{i}^{*})_{i=1}^{m}$ satisfies $\theta_{1}^{*}\geq 1$. Then Corollary 25 implies that the algorithm $\mathcal{A}(F,(\theta^{*})^{\nicefrac{{1}}{{n}}})$ is $\alpha$-consistent and $\beta$-robust. The results are plotted in Figure 1(a).
Intuitively, as $m\to\infty$, the step function approximation tightens. In experiments, we set $m=300$. For details, see the implementation of the backward recursion as well as the solved $\theta_{i}^{*}$’s in our code repository.
Figure 3 illustrates the case when $\beta=0.01$ and $m=300$ yields $\alpha\approx 0.6908$ with $\theta^{*}_{1}\approx 1.0165$. This implies that $\mathcal{A}(F,(\theta^{*})^{\nicefrac{{1}}{{n}}})$ with $\theta^{*}$ being the step function with step values $(\min\{\theta_{i}^{*},1\})_{i=1}^{300}$ is an $\alpha\approx 0.6908$-consistent and $\beta=0.01$-robust algorithm.
Figure 3: Example solution of $\theta$ when $\beta=0.01$, $m=300$, and $\alpha\approx 0.6908$.
Appendix C Missing Proofs for MaxProb Hardness
C.1 Minor-Oblivious Algorithms: Proof of Proposition 17
The key observation is that, when the current value is best-so-far and equal across sequences, the conditional probability of stopping at time $t$ is invariant across all truncation levels $k\geq\ell$ for which the conditioning event has positive probability. This is formalized in the following lemma.
Lemma 26.
For any $t\in[n]$ and any $\ell\in[K]$, the conditional probabilities
$\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}=t\mid\textsc{Alg}\geq t,\ x_{t}=\max\bm{x}_{1:t}=\ell\,\big]$
are identical for all $k$ with $\ell\leq k\leq K$.
Proof.
By Bayes’ rule, we can obtain that
$\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}=t\mid\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]~=~\frac{\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}=t\,,\,x_{t}=\ell\mid\max\bm{x}_{1:t}=\ell\,\big]}{\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}\geq t\,,\,x_{t}=\ell\mid\max\bm{x}_{1:t}=\ell\,\big]}~.$
Denote by $S\subseteq[K]^{t}$ the set of all $\bm{x}_{1:t}$ with $x_{t}=\max\bm{x}_{1:t}=\ell$. The numerator can be written as
$\sum_{\bm{x}_{1:t}\in S}\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}=t\mid\bm{x}_{1:t}\,\big]\cdot\operatorname{\mathbf{Pr}}_{k}\big[\,\bm{x}_{1:t}\mid\max\bm{x}_{1:t}=\ell\,\big]~,$
where the second factor is the conditional probability of realizing $\bm{x}_{1:t}$ given $\max\bm{x}_{1:t}=\ell$. This is invariant for all $k$ with $\ell\leq k\leq K$ by Lemma 16. The same holds for the denominator. Hence, the ratio is invariant. ∎
Given any algorithm $\mathcal{A}$, whose index of accepted value is Alg, we construct a minor-oblivious algorithm $\mathcal{A}^{\prime}$ with conditional acceptance probabilities
$r^{\prime}_{t}(\ell)~\coloneqq~\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}=t\mid\textsc{Alg}\geq t,\ x_{t}=\max\bm{x}_{1:t}=\ell\,\big]~.$
That is, the probability of accepting a best-so-far value $x_{t}=\ell$ equals the probability that $\mathcal{A}$ accepts at time $t$, conditioned on 1) $x_{t}=\ell$; 2) $x_{t}$ is best-so-far; and 3) $\mathcal{A}$ rejects the first t1t-1 values. By Lemma 26, we may define this conditional probability with respect to prior $F_{K}$ without loss of generality.
Finally, we show that algorithm $\mathcal{A}^{\prime}$ accepts the maximum value with the same probability as $\mathcal{A}$. Let $\textsc{Alg}^{\prime}$ be the index of the accepted value by algorithm $\mathcal{A}^{\prime}$. For any prior $F_{k}$, observe that
$\operatorname{\mathbf{Pr}}_{k}\left[x_{\textsc{Alg}}=x_{\textsc{Opt}}\right]$ $~=~\sum_{t\in[n]}\sum_{\ell\in[k]}\,\operatorname{\mathbf{Pr}}_{k}\left[\max\bm{x}_{t+1:n}\leq\ell\right]\cdot\operatorname{\mathbf{Pr}}\nolimits_{k}\left[\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\right]$
$\qquad\qquad\qquad\cdot\operatorname{\mathbf{Pr}}\left[\textsc{Alg}=t\mid\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\right]~.$
Similarly, we have
$\operatorname{\mathbf{Pr}}_{k}\left[x_{\textsc{Alg}^{\prime}}=x_{\textsc{Opt}}\right]$ $~=~\sum_{t\in[n]}\sum_{\ell\in[k]}\,\operatorname{\mathbf{Pr}}_{k}\left[\max\bm{x}_{t+1:n}\leq\ell\right]\cdot\operatorname{\mathbf{Pr}}\nolimits_{k}\left[\textsc{Alg}^{\prime}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\right]$
$\qquad\qquad\qquad\cdot\operatorname{\mathbf{Pr}}\left[\textsc{Alg}^{\prime}=t\mid\textsc{Alg}^{\prime}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\right]~.$
Comparing the two expressions, the first parts coincide, and the third parts are equal by the construction of $\mathcal{A}^{\prime}$. It remains to prove that for any $t\in[n]$ and any $k,\ell\in[K]$:
$\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}^{\prime}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]~=~\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]~.$ (22)
We next prove this by a joint induction with the following claim:
$\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}^{\prime}\geq t\,,\,x_{t}<\max\bm{x}_{1:t}=\ell\,\big]~=~\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}\geq t\,,\,x_{t}<\max\bm{x}_{1:t}=\ell\,\big]~.$ (23)
The base case of $t=1$ holds trivially, as the probabilities are independent of the algorithms’ decisions. Next, suppose that both equalities hold for some $t\geq 1$. We will prove them for $t+1$. Enumerating over the state at step $t$, we have:
$\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}\geq t+1\,,\,x_{t+1}=\max\bm{x}_{1:t+1}=\ell\,\big]$
$\quad~=~\sum_{m\in[\ell]}\operatorname{\mathbf{Pr}}_{k}\big[\,\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=m\,\big]\cdot\operatorname{\mathbf{Pr}}_{k}\big[x_{t+1}=\ell\big]$ (24)
$\qquad\qquad\cdot\,\big(1-\operatorname{\mathbf{Pr}}[\textsc{Alg}=t\mid\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=m]\big)$ (25)
$\quad\qquad+\sum_{m\in[\ell]}\operatorname{\mathbf{Pr}}\nolimits_{k}\big[\textsc{Alg}\geq t\,,\,x_{t}<\max\bm{x}_{1:t}=m\big]\cdot\operatorname{\mathbf{Pr}}_{k}\big[x_{t+1}=\ell\big]\,.$ (26)
By the induction hypothesis, we can replace Alg with $\textsc{Alg}^{\prime}$ in (24) and (26) while preserving the value. Further, we can also replace Alg with $\textsc{Alg}^{\prime}$ in (25) by the construction of algorithm $\mathcal{A}^{\prime}$. Hence, Equation 22 holds for $t+1$ as well.
The proof of Equation 23 for $t+1$ is almost verbatim.
C.2 Relating Rejection and Acceptance Probabilities: Proof of Lemma 18
From Rejection Probabilities to Acceptance Probabilities.
By definition, we have:
$\textsc{Acc}_{t}(\ell)~=$~ $\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}=t\mid\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]$
$~=$~ $1-\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}>t\mid\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]\,.$
By Bayes’ rule, the probability on the right-hand side can be written as
$\operatorname{\mathbf{Pr}}\big[\,\textsc{Alg}>t\mid\textsc{Alg}\geq t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]~=~\frac{\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}>t\,,\,x_{t}=\max\bm{x}_{1:t}=\ell\,\big]}{\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}\geq t\,,\,x_{t}=\max\mathbf{x}_{1:t}=\ell\,\big]}\,.$
Here we consider the prior $F_{K}$ without loss of generality, because using any other $F_{k}$ multiplies numerator and denominator by the same $\tilde{F}(k)^{-t}$ factor according to Definition 7.
The numerator further equals
$\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}>t\,,\,\max\bm{x}_{1:t}=\ell\,\big]~-~\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}>t\,,\,x_{t}<\max\bm{x}_{1:t}=\ell\,\big]$
$\quad~=~\operatorname{\mathbf{Pr}}_{K}\big[\,\max\bm{x}_{1:t}=\ell\,\big]\cdot\textsc{Rej}_{t}(\ell)-\operatorname{\mathbf{Pr}}_{K}\big[\,x_{t}<\max\bm{x}_{1:t-1}=\ell\,\big]\cdot\textsc{Rej}_{t-1}(\ell)~.$
The denominator equals
$\sum_{m\in[\ell]}\operatorname{\mathbf{Pr}}_{K}\big[\,\textsc{Alg}\geq t\,,\,\max\bm{x}_{1:t-1}=m\,,\,x_{t}=\ell\,\big]=\sum_{m\in[\ell]}\operatorname{\mathbf{Pr}}_{K}\big[\,\max\bm{x}_{1:t-1}=m\,,\,x_{t}=\ell\,\big]\cdot\textsc{Rej}_{t-1}(m)\,.$
Finally, putting these together proves Lemma 18.
From Acceptance Probabilities to Rejection Probabilities.
We remark that the relation between acceptance probabilities and rejection probabilities is one-to-one. Given the acceptance probabilities, we can inductively deduce the rejection probabilities using Lemma 18. Restating Equation 12 from Lemma 18:
$\textsc{Acc}_{t}(\ell)~=~1-\frac{\textsc{Rej}_{t}(\ell)\cdot\operatorname{\mathbf{Pr}}_{K}\big[\max\bm{x}_{1:t}=\ell\,\big]-\textsc{Rej}_{t-1}(\ell)\cdot\operatorname{\mathbf{Pr}}_{K}\big[\,x_{t}<\max\bm{x}_{1:t-1}=\ell\,\big]}{\sum_{m\in[\ell]}\textsc{Rej}_{t-1}(m)\cdot\operatorname{\mathbf{Pr}}_{K}\big[\max\bm{x}_{1:t-1}=m,x_{t}=\ell\,\big]}\,.$
Given $\textsc{Rej}_{t-1}(\cdot)$ and $\textsc{Acc}_{t}(\cdot)$ for any $t\geq 1$, we can derive $\textsc{Rej}_{t}(\cdot)$ from the equation above.